### v1

# Global Behavior of Learning Dynamics in Zero-Sum Games with Memory Asymmetry

Abstract
This study examines the global behavior of dynamics in learning in games between two players, X and Y. We consider the simplest situation for memory asymmetry between two players: X memorizes the other Y’s previous action and uses reactive strategies, while Y has no memory. Although this memory complicates the learning dynamics, we discover two novel quantities that characterize the global behavior of such complex dynamics. One is an extended Kullback-Leibler divergence from the Nash equilibrium, a well-known conserved quantity from previous studies. The other is a family of Lyapunov functions of X’s reactive strategy. These two quantities capture the global behavior in which X’s strategy becomes more exploitative, and the exploited Y’s strategy converges to the Nash equilibrium. Indeed, we theoretically prove that Y’s strategy globally converges to the Nash equilibrium in the simplest game equipped with an equilibrium in the interior of strategy spaces. Furthermore, our experiments also suggest that this global convergence is universal for more advanced zero-sum games than the simplest game. This study provides a novel characterization of the global behavior of learning in games through a couple of indicators.
1 Introduction
Learning in games targets how multiple agents learn their optimal strategies in the repetition of games [1]. The set of such players’ best strategies is defined as Nash equilibrium [2], where every player has no motivation to change his/her strategy. However, this equilibrium is hard to compute in general because one’s best strategy depends on the others’ strategies. Indeed, the behavior of multi-agent learning is complicated in zero-sum games, where players conflict in their payoffs. Even when players try to learn their optimal strategies there, their strategies often cycle around the Nash equilibrium and fail to converge to the equilibrium.
In order to understand such strange behaviors, which are unique in multi-agent learning, the dynamics of how multiple agents learn their strategies, say, learning dynamics, are frequently studied [3, 4, 5, 6]. The representative dynamics of interest are the replicator dynamics, which is based on the evolutionary dynamics in biology [7, 8, 9, 10, 11]. These dynamics are also known as the multiplicative weight updates (MWU) in its discrete-time version [12, 13]. Furthermore, their connection to other representative learning dynamics, such as gradient ascent [14, 15, 16, 17] and Q-learning [18, 19, 20], should be noted. Such replicator dynamics are known to be characterized by Kullbuck-Leibler (KL) divergence, which is the distance from the Nash equilibrium to the players’ present strategies. This KL divergence is conserved during the learning dynamics, and the distance from the Nash equilibrium is invariant. Follow the Regularized Leader (FTRL) is a class of learning algorithms including the replicator dynamics and also has its conserved quantity, which is the summation of divergences for all the players [21, 22]. To summarize, such complex learning dynamics have been discussed based on their conserved quantity.
In this study, we define memory as an agent’s ability to change its action choice depending on the outcome of past games. By definition, this memory allows the agent to make more complex and intelligent decisions. When memory is introduced into a normal-form game, the players can achieve a wider range of strategies as the Nash equilibria (known as Folk theorem [23]). Furthermore, memory is also introduced into learning algorithms, such as replicator dynamics [24, 25, 26, 27] and Q-learning [28, 29, 30, 31, 32]. Here, since this memory causes feedback from the past, the global dynamics of such learning algorithms become more complex. Indeed, replicator dynamics diverge from the Nash equilibrium under symmetric memory lengths between players [26], while converging under asymmetric memory lengths [27]. Here, KL divergence is no longer useful to capture the global dynamics because it increases or decreases over time. The analysis of the dynamics in with-memory games is limited to the local, linearized stability analysis in the vicinity of Nash equilibrium [27]. To summarize, since memory crucially complicates learning dynamics, the global behavior of the dynamics is still unexplored.
This study provides the first theoretical analysis of the global behavior of learning in with-memory games. We assume games where their memory structure is simplest and asymmetric; One side adopts a reactive strategy that can memorize the other’s previous action [33, 34, 35, 24, 36, 25], while the other has no memory. In order to characterize the global behavior of such with-memory games, we extend KL divergence and prove that such extended divergence increases or decreases with time depending on the reactive strategy (see Fig. 1A). Furthermore, we propose a family of Lyapunov functions that characterize the dynamics of reactive strategy (see Fig. 1B). These Lyapunov functions show that the with-memory side monotonically learns to exploit the no-memory side. As an application of these functions, we prove that the convergence from arbitrary initial strategies, i.e., global convergence, occurs in matching pennies. We also experimentally confirm that such global convergence is observed in general zero-sum games.
Figure 1: A. Illustration of the global behavior of the conditional divergence, $D(\bm{X},\bm{y})$. Three trajectories (red, black, and blue) are plotted with the Nash equilibrium (the black star marker). The horizontal and vertical axes show X’s strategy ($x_{1}^{{\rm st}}$) and Y’s strategy ($y_{1}$) in matching pennies (formulated in Fig. 2). This divergence decreases (red: $\dot{D}<0$), cycles (black: $\dot{D}=0$), or increases (blue: $\dot{D}>0$) with time. These three lines are plotted for the different initial strategies, i.e., $\bm{X}$ and $\bm{y}$. B. Illustration of the global behavior of the family of Lyapunov functions, $H(\bm{X};\bm{\delta})$. The colored line shows a trajectory (from purple to red) of Lyapunov functions $H_{1}$, $H_{2}$, and $H_{3}$, each of which is $H(\bm{X};\bm{\delta})$ for some specific $\bm{\delta}$. The gray broken lines are the projections of the black solid line to $H_{1}$-$H_{2}$, $H_{2}$-$H_{3}$, and $H_{3}$-$H_{1}$ planes. All of $H_{1}$, $H_{2}$, and $H_{3}$ monotonically increase with time.
2 Preliminary
2.1 Settings
First, we formulate two-player normal-form games. We consider two players, denoted as X and Y. X’s actions are denoted as $\{a_{i}\}_{1\leq i\leq m_{\sf X}}$, while Y’s are $\{b_{j}\}_{1\leq j\leq m_{\sf Y}}$. When X and Y choose $a_{i}$ and $b_{j}$, they obtain the payoffs of $u_{ij}\in\mathbb{R}$ and $v_{ij}\in\mathbb{R}$, respectively. Thus, all their possible payoffs are given by the matrices, $\bm{U}:=(u_{ij})_{ij}$ and $\bm{V}:=(v_{ij})_{ij}$. Here, when $\bm{V}=-\bm{U}$ holds, the games are called zero-sum. Although our formulation of learning algorithms can be used for general games, this study focuses on zero-sum games.
Figure 2: Illustration of games between reactive and zero-memory strategies. The area surrounded by the magenta dotted line shows the normal-form game. In each round, X chooses action $i=1$ or $2$ in the row, following its strategy, i.e., the probability distribution of $\bm{x}=(x_{1},x_{2})$. On the other hand, Y chooses action $j=1$ or $2$ in the column, following its strategy, i.e., the probability distribution of $\bm{y}=(y_{1},y_{2})$. Depending on their choices $i$ and $j$, X receives a payoff $u_{ij}$, given by a matrix form of $\bm{U}=(u_{ij})_{i,j}=((u_{11},u_{12}),(u_{21},u_{22}))$. Furthermore, in zero-sum games, Y receives $-u_{ij}$. Especially in the matching pennies, their actions of $1$ ($2$) correspond to the choice of “head” (“tail”) of a coin. When their choices match $i=j$, X wins, i.e., $u_{11}=u_{22}=1$ (the orange blocks). Else when their choices mismatch $i\neq j$, Y wins, i.e., $u_{12}=u_{21}=-1$ (the blue blocks). The area outside of the magenta dotted line shows the difference due to an effect of memory. The gray box shows that X memorizes Y’s previous action, represented as $j=1$ or $2$. Thus, X uses a reactive strategy and can choose its action with the conditional probability of $x_{1|j}$ and $x_{2|j}$ for Y’s previous action.
We assume that X can use reactive strategies, i.e., can change its action choice depending on the other’s previous action. This reactive strategy is denoted as $\bm{X}:=(x_{i|j})_{1\leq i\leq m_{\sf X},1\leq j\leq m_{\sf Y}}\in\prod_{1\leq j\leq m_{\sf Y}}\Delta^{m_{\sf X}-1}$, a matrix composed of $m_{\sf Y}$ vectors each of which are an element of a $m_{\sf X}-1$-dimensional simplex. Here, $x_{i|j}$ means the probability that X chooses $a_{i}$ in the condition when Y’s previous action is $b_{j}$. Thus, $\Sigma_{i}x_{i|j}=1$ should be satisfied for all $j$. On the other hand, Y only can use classical mixed strategies and choose its own action without reference to the previous actions. This mixed strategy is denoted as $\bm{y}=(y_{j})_{1\leq j\leq m_{\sf Y}}\in\Delta^{m_{\sf Y}-1}$, a vector which is an element of a $(m_{\sf Y}-1)$-dimensional simplex. Thus, $\Sigma_{j}y_{j}=1$ should be satisfied.
2.2 Stationary state and expected payoff
We now discuss the stationary state and expected payoff of repeated games. Since Y determines its action independent of the outcomes of previous rounds, X’s stationary strategy, defined as $\bm{x}^{{\rm st}}:=(x_{i}^{{\rm st}})_{1\leq i\leq m_{\sf X}}$, is given by $x_{i}^{{\rm st}}(\bm{x}_{i},\bm{y})=\Sigma_{j}x_{i|j}y_{j}$. Here, $x_{i}^{{\rm st}}$ means the probability that X chooses $a_{i}$ in the stationary state. Furthermore, the stationary state is described as $\bm{P}^{{\rm st}}:=\bm{x}^{{\rm st}}\otimes\bm{y}$ with use of $\bm{x}^{{\rm st}}$ and $\bm{y}$. Last, X’s expected payoff is given by $u^{{\rm st}}(\bm{x}^{{\rm st}},\bm{y}):=\Sigma_{i}\Sigma_{j}u_{ij}p_{ij}^{{\rm st}}=\Sigma_{i}\Sigma_{j}u_{ij}x_{i}^{{\rm st}}y_{j}$.
2.3 Nash equilibrium
We now define the Nash equilibrium in the normal-form game. Here, note that this equilibrium is based on games without memories, where X’s strategy does not refer to the past games, i.e., $\bm{x}:=(x_{i})_{i}$. By using the expected payoff $u^{{\rm st}}(\bm{x},\bm{y})$ for games without memories, the Nash equilibrium $(\bm{x}^{*},\bm{y}^{*})$ is formulated as
$\begin{cases}\bm{x}^{*}={\rm argmax}_{\bm{x}}u^{{\rm st}}(\bm{x},\bm{y}^{*})\\ \bm{y}^{*}={\rm argmin}_{\bm{y}}u^{{\rm st}}(\bm{x}^{*},\bm{y})\\ \end{cases}.$ (1)
From the definition, $u^{{\rm st}}(\bm{x},\bm{y})$ is the linear function for $\bm{x}$ and $\bm{y}$, and the Nash equilibrium condition is characterized by the gradient of such expected payoffs as
$\begin{cases}\partial u^{{\rm st}}/\partial x_{i}=\Sigma_{j}u_{ij}y_{j}^{*}=C&(x_{i}^{*}>0)\\ \partial u^{{\rm st}}/\partial x_{i}=\Sigma_{j}u_{ij}y_{j}^{*}<C&(x_{i}^{*}=0)\\ \end{cases},$ (2)
$\begin{cases}\partial u^{{\rm st}}/\partial y_{j}=\Sigma_{j}u_{ij}x_{i}^{*}=C&(y_{i}^{*}>0)\\ \partial u^{{\rm st}}/\partial y_{i}=\Sigma_{j}u_{ij}x_{i}^{*}>C&(y_{i}^{*}=0)\\ \end{cases}.$ (3)
From these conditions, for all $i$ and $j$ such that $x_{i}^{*}>0$ and $y_{j}^{*}>0$ hold, respectively, we obtain
$\Sigma_{i}u_{ij}x_{i}^{*}=\Sigma_{j}u_{ij}y_{j}^{*}=:u^{*}.$ (4)
Let us interpret this equation. First, $\Sigma_{i}u_{ij}x_{i}^{*}=u^{*}$ means that when X takes its Nash equilibrium strategy, its own payoff is fixed to $u^{*}$, independent of Y’s strategy. On the other hand, $\Sigma_{j}u_{ij}y_{j}^{*}=u^{*}$ similarly means that Y’s Nash equilibrium strategy fixes X’s payoff to $u^{*}$. In other words, either X or Y takes its Nash equilibrium strategy, their payoffs are fixed. This is the special property in zero-sum games.
2.4 Learning algorithm: replicator dynamics
Let us define the replicator dynamics as a representative learning algorithm. X’s and Y’s replicator dynamics are formulated as
$\dot{x}_{i|j}$ $=x_{i|j}\left(\frac{{\rm d}u^{{\rm st}}}{{\rm d}x_{i|j}}-\Sigma_{i}x_{i|j}\frac{{\rm d}u^{{\rm st}}}{{\rm d}x_{i|j}}\right),$ (5)
$\dot{y}_{j}$ $=-y_{j}\left(\frac{{\rm d}u^{{\rm st}}}{{\rm d}y_{j}}-\Sigma_{j}y_{j}\frac{{\rm d}u^{{\rm st}}}{{\rm d}y_{j}}\right).$ (6)
Here, following the theorems in [26], X’s and Y’s replicator dynamics include the gradient for the expected payoff $u^{{\rm st}}$. Thus, the update of X’s strategy increases its payoff $u^{{\rm st}}$, while that of Y’s strategy decreases the other’s payoff $u^{{\rm st}}$.
3 Novel quantities characterizing learning dynamics
This section analyzes the dynamics of Eqs. (5) and (6). First, we compute in detail the gradient terms, which appear to be complex. Next, as a preliminary, we define positive definite matrices for some special vectors. Based on this definition, we introduce two quantities characterizing the dynamics of Eqs. (5) and (6): An extended KL divergence and a family of Lyapunov functions.
3.1 Polynomial expressions of learning dynamics
First, the gradient terms in Eqs. (5) and (6) are computed as
$\frac{{\rm d}u^{{\rm st}}(\bm{x}^{{\rm st}}(\bm{X},\bm{y}),\bm{y})}{{\rm d}x_{i|j}}$ $=\frac{\partial x_{i}^{{\rm st}}(\bm{x}_{i},\bm{y})}{\partial x_{i|j}}\frac{\partial u^{{\rm st}}(\bm{x}^{{\rm st}},\bm{y})}{\partial x_{i}^{{\rm st}}}$ (7)
$=y_{j}\Sigma_{j^{\prime}}u_{ij^{\prime}}y_{j^{\prime}},$ (8)
$\frac{{\rm d}u^{{\rm st}}(\bm{x}^{{\rm st}}(\bm{X},\bm{y}),\bm{y})}{{\rm d}y_{j}}$ $=\frac{\partial u^{{\rm st}}(\bm{x}^{{\rm st}},\bm{y})}{\partial y_{j}}+\Sigma_{i}\frac{\partial x_{i}^{{\rm st}}(\bm{x}_{i},\bm{y})}{\partial y_{j}}\frac{\partial u^{{\rm st}}(\bm{x}^{{\rm st}},\bm{y})}{\partial x_{i}^{{\rm st}}}$ (9)
$=\Sigma_{i}u_{ij}x_{i}^{{\rm st}}+\Sigma_{i}x_{i|j}\Sigma_{j^{\prime}}u_{ij^{\prime}}y_{j^{\prime}}.$ (10)
Here, we remark that Eqs. (5) and (6) are nonlinear functions of $\bm{X}$ and $\bm{y}$, which is a feature of learning in with-memory games. Notably, however, these equations are polynomial expressions with $\bm{X}$ and $\bm{y}$. Such polynomial expressions cannot be seen in the games of other memory lengths [26, 27] but are special in the games between reactive and no memory strategies.
3.2 Positive definiteness for zero-sum vectors
Next, let us introduce a definiteness of matrices. Here, however, this definite matrix is for vectors whose elements are summed to 00, named “zero-sum vectors”. In mathematics, zero-sum vector $\bm{\delta}:=(\delta_{k})_{k}$ satisfies $\Sigma_{k}\delta_{k}=0$ but $\bm{\delta}\neq\bm{0}$.
Definition 1 (Positive definiteness for zero-sum vectors).
A square matrix $\bm{M}$ is “positive definite for zero-sum vectors” when for all vectors $\bm{\delta}\neq\bm{0}$ such that $\Sigma_{k}\delta_{k}=0$, $\bm{\delta}\cdot(\bm{M}\bm{\delta})<0$ holds.
The positive definiteness for zero-sum vectors connects with an ordinary positive definiteness by a simple transformation of a matrix (see Appendix B for details).
3.3 Extended Kullback-Leibler divergence
The first quantity is an extended version of divergence. Before considering the extension, we introduce the classical version of divergence $D_{\rm c}$, which is the function of X’s mixed strategies ($\bm{x}:=(x_{i})_{1\leq i\leq m_{\sf X}}\in\Delta^{m_{\sf X}-1}$) and Y’s mixed strategies ($\bm{y}$) as
$D_{\rm c}(\bm{x},\bm{y})$ $:=D_{\rm KL}(\bm{x}^{*}\|\bm{x})+D_{\rm KL}(\bm{y}^{*}\|\bm{y}),$ (11)
$D_{\rm KL}(\bm{p}^{*}\|\bm{p})$ $:=\bm{p}^{*}\cdot\log\bm{p}^{*}-\bm{p}^{*}\cdot\log\bm{p}.$ (12)
We now give an intuitive interpretation of this quantity. First, $D_{\rm KL}(\bm{p}^{*}\|\bm{p})$ is the KL divergence, meaning the distance from the reference point $\bm{p}^{*}$ to the target point $\bm{p}$. Thus, $D_{\rm c}(\bm{x},\bm{y})$ means the total distance from the Nash equilibrium $(\bm{x}^{*},\bm{y}^{*})$ to the current state $(\bm{x},\bm{y})$.
Let us extend the classical divergence to the case of this study, where X refers to the previous action of the other and can use reactive strategies $\bm{X}$. This extended divergence, i.e., $D(\bm{X},\bm{y})$, is named the “conditional-sum” divergence, formulated as
$D(\bm{X},\bm{y})$ $:=\Sigma_{j}D_{\rm KL}(\bm{x}^{*}\|\bm{x}_{j})+D_{\rm KL}(\bm{y}^{*}\|\bm{y}).$ (13)
We now remark the difference between $D(\bm{X},\bm{y})$ and $D_{\rm c}(\bm{x},\bm{y})$. Recall that X’s reactive strategy is defined as $(\bm{x}_{j})_{1\leq j\leq m_{\sf Y}}$, which shows how to choose its action with the condition that Y chose $b_{j}$ in the previous round. Hence, $D(\bm{X},\bm{y})$ represents the summation of KL divergence from $\bm{x}^{*}$ to $\bm{x}_{j}$ for all the conditions of $j$. Here, we also remark that when the reactive strategy does not use memory, i.e., $\bm{x}_{j}=\bm{x}$ for all $j$, this conditional-sum divergence also captures the behavior of the classical divergence (see Appendix C for details).
This conditional-sum divergence satisfies the following theorem (see Appendix A.1 for its full proof).
Theorem 1 (Monotonic decrease of $D$ for positive definite $\bm{X}^{\rm T}\bm{U}$).
If $\bm{X}^{\rm T}\bm{U}$ is positive definite for zero-sum vector, $D^{\dagger}(\bm{X};{\rm d}\bm{y}):=\dot{D}(\bm{X},\bm{y})<0$ for all ${\rm d}\bm{y}\neq\bm{0}$.
Proof Sketch. We calculation $\dot{D}(\bm{X},\bm{y})$ in practice. In the calculation, the contribution of X’s gradient (Eq. (8)) cancels out the contribution of the first term of the gradient of Y (Eq. (10)). (Here, we remark that the same canceling out also occurs in the calculation for the conservation of the classical divergence $D_{\rm c}(\bm{x},\bm{y})$ in games without memory.) However, the contribution of the second term of Eq. (10) is special in games of a reactive strategy. By using the constant payoff condition in the Nash equilibrium (Eqs. (4)), we obtain
$\dot{D}(\bm{X},\bm{y})=-{\rm d}\bm{y}^{\rm T}\bm{X}^{\rm T}\bm{U}{\rm d}\bm{y}\ (=:D^{\dagger}(\bm{X};{\rm d}\bm{y})),$ (14)
where we defined ${\rm d}\bm{y}:=\bm{y}-\bm{y}^{*}$, which means the difference from Y’s equilibrium strategy and is a zero-sum vector. Thus, when $\bm{X}^{\rm T}\bm{U}$ is positive definite for zero-sum vectors, ${\rm d}\bm{y}^{\rm T}\bm{X}^{\rm T}\bm{U}{\rm d}\bm{y}$ is always positive, leading to $D^{\dagger}(\bm{X};{\rm d}\bm{y})<0$ for all ${\rm d}\bm{y}\neq\bm{0}$. ∎
If $\bm{X}^{\rm T}\bm{U}$ are always positive definite for zero-sum vectors, Thm. 1 gives the global behavior of learning dynamics as below.
Corollary 1 (Convergence of Y’s strategy).
If $\bm{X}^{\rm T}\bm{U}$ continues to be positive definite for zero-sum vectors, $\bm{y}=\bm{y}^{*}$ holds after sufficiently long time passes.
Proof.
This corollary is straightforwardly proved by Thm. 1. If $\bm{X}^{\rm T}\bm{U}$ is positive definite and if $\bm{y}\neq\bm{y}^{*}$, $D^{\dagger}(\bm{X};{\rm d}\bm{y})<0$ holds. Here, $D(\bm{X},\bm{y})$ has its lower bound ($\min_{\bm{X},\bm{y}}D(\bm{X},\bm{y})=0$). Thus, $D^{\dagger}(\bm{X};{\rm d}\bm{y})=0\Leftrightarrow\bm{y}=\bm{y}^{*}$ should be satisfied eventually. ∎
3.4 Family of Lyapunov functions
Furthermore, we introduce a Lyapunov function, which characterizes the learning dynamics of X’s reactive strategy. Based on an arbitrary zero-sum vector $\bm{\delta}:=(\delta_{i})_{1\leq i\leq m_{\sf X}}$, this function is defined as
$H(\bm{X};\bm{\delta}):=\bm{\delta}^{\rm T}\bm{U}\log\bm{X}^{\rm T}\bm{\delta}.$ (15)
The following theorem holds for this Lyapunov function (see Appendix A.2 for its full proof).
Theorem 2 (Monotonic increase of $H$).
For all $\bm{\delta}$ such that $\Sigma_{i}\delta_{i}=0$, $H(\bm{X};\bm{\delta})$ monotonically increases with time.
Proof Sketch. By using Eq. (4), we calculate
$\dot{H}(\bm{X};\bm{\delta})=(\bm{\delta}^{\rm T}\bm{U}{\rm d}\bm{y})^{2}\ (=:H^{\dagger}(\bm{y};\bm{\delta})).$ (16)
This means that $H^{\dagger}(\bm{y};\bm{\delta})\geq 0$ for all $\bm{\delta}$. ∎
Let us interpret the function of $H(\bm{X};\bm{\delta})$. For simplicity, we consider a simple class of zero-sum vectors, $\bm{\delta}=\bm{e}_{i}-\bm{e}_{i^{\prime}}$ for some $i$ and $i^{\prime}\neq i$, where $\bm{e}_{i}$ denotes the unit vector for $i$-th element. Then, we calculate the function as $H(\bm{X};\bm{e}_{i}-\bm{e}_{i^{\prime}})=\bm{u}_{i}\cdot\log\bm{x}_{i}-\bm{u}_{i}\cdot\log\bm{x}_{i^{\prime}}-\bm{u}_{i^{\prime}}\cdot\log\bm{x}_{i}+\bm{u}_{i^{\prime}}\cdot\log\bm{x}_{i^{\prime}}$. The monotonic increase of $H(\bm{X};\bm{e}_{i}-\bm{e}_{i^{\prime}})$ shows that $\log\bm{x}_{i}$ becomes correlated to $\bm{u}_{i}$, and $\log\bm{x}_{i^{\prime}}$ to $\bm{u}_{i^{\prime}}$. Here, $\bm{x}_{i}=(x_{i|j})_{j}$ represents the vector of the probabilities to choose $i$-th action in the condition that the other chose $j$-th action in the previous round. $\bm{u}_{i}=(u_{ij})_{j}$ also represents the vector of the payoffs for action $i$. Thus, the correlation between $\log\bm{x}_{i}$ and $\bm{u}_{i}$ means that the larger its reward $u_{ij}$ is, the more frequently X returns its $i$-th action against the other’s $j$-th action, in other words, X becomes more exploitative.
This theorem characterizes the state of learning dynamics after a sufficiently long time passes, as follows.
Corollary 2 (Convergence of Y’s strategy or divergence of $H$).
After sufficiently long time passes, $\bm{y}=\bm{y}^{*}$ holds or $H(\bm{X};\bm{\delta})$ diverges to infinity for all $\bm{\delta}\neq\bm{0}$ such that $\Sigma_{i}\delta_{i}=0$.
Proof.
Since $H(\bm{X};\bm{\delta})\geq 0$ always holds for all $\bm{\delta}$, either $H(\bm{X};\bm{\delta})=0$ or $H(\bm{X};\bm{\delta})>0$ holds in the final state. The former is equivalent to $\bm{y}=\bm{y}^{*}$, while the latter is equivalent to $H(\bm{X};\bm{\delta})\to\infty$. ∎
4 Global behavior of learning dynamics
Now, Cors. 1 and 2 are utilized to capture the global behavior of the learning dynamics. One of the global behaviors seen in the zero-sum games is the global convergence to the Nash equilibrium independent of X’s and Y’s initial strategies. First, Cor. 1 guarantees that after a sufficiently long time passes, either Y’s strategy converges to the equilibrium or $H(\bm{X};\bm{\delta})$ diverges. Since the former trivially leads to such global convergence, we now consider the latter condition, i.e., $H(\bm{X};\bm{\delta})=\bm{\delta}^{\rm T}\bm{U}\log\bm{X}^{\rm T}\bm{\delta}\to\infty$. Here, $H(\bm{X};\bm{\delta})=\bm{\delta}^{\rm T}\bm{U}\log\bm{X}^{\rm T}\bm{\delta}\to\infty$ naively connects to $-D^{\dagger}(\bm{X},{\rm d}\bm{y})={\rm d}\bm{y}^{\rm T}\bm{X}^{\rm T}\bm{U}{\rm d}\bm{y}>0$ (explained in the next paragraph), meaning that $\bm{X}^{\rm T}\bm{U}$ is positive definite for zero-sum vectors. If so, Cor. 2 becomes applicable, and either Y’s strategy converges to the Nash equilibrium.
We explain the connection between $H(\bm{X};\bm{\delta})=\bm{\delta}^{\rm T}\bm{U}\log\bm{X}^{\rm T}$ and $-D^{\dagger}(\bm{X},{\rm d}\bm{y})={\rm d}\bm{y}^{\rm T}\bm{X}^{\rm T}\bm{U}{\rm d}\bm{y}>0$. Here, $H(\bm{X};\bm{\delta})$ and $-D^{\dagger}(\bm{X},{\rm d}\bm{y})$ have two commonalities. The first one is that $\bm{\delta}$ and ${\rm d}\bm{y}$ are both zero-sum vectors in common. The other one is that both $H$ and $-D^{\dagger}$ are given by the product of X’s strategy $\bm{X}$ and the payoff matrix $\bm{U}$. This means that the more $\bm{X}$ is correlated to $\bm{U}$, the larger both $H$ and $-D^{\dagger}$ are. Here, however, we also remark two differences of $H$ from $-D^{\dagger}$; $H$ considers the logarithm in X’s strategy $\bm{X}$ and the reversed order in the product of $\bm{X}$ and $\bm{U}$.
In the following, we prove such global convergence of either player’s strategy in the matching pennies, the simplest game equipped with a full-support equilibrium. We also experimentally show global convergence in more complicated games than the matching pennies later.
4.1 Example: Matching-pennies
Let us define the matching-pennies game (see Fig. 2 for the illustration of its payoff matrix). This game considers the action numbers of $m_{\sf X}=m_{\sf Y}=2$ and the payoff matrix of $\bm{U}=((u_{11},u_{12}),(u_{21},u_{22}))=((+1,-1),(-1,+1))$. The Nash equilibrium of this game is only $\bm{x}^{*}=\bm{y}^{*}=(1/2,1/2)$. This game is the simplest example of games that have a full-support equilibrium. In addition, it has been known that the replicator dynamics in games without memories cycle around the Nash equilibrium and cannot reach the equilibrium. Nevertheless, the learning dynamics in games of a reactive strategy achieve the convergence to the Nash equilibrium for Y’s strategy. The following theorem holds (see Appendix A.3 for its full proof).
Theorem 3 (Global convergence in matching-pennies).
In the matching-pennies ($\bm{U}=((+1,-1),(-1,+1))$), Y’s strategy converges to the Nash equilibrium ($\bm{y}=\bm{y}^{*}$), independent of both the players’ initial strategies.
Proof Sketch. First, we define a special notation in two-action games; $(x_{1|j},x_{2|j})=:(x_{j},1-x_{j})$ for $j=1,2$. We also define the logit function of $x_{j}$ as $q_{j}:=\log x_{j}-\log(1-x_{j})$. Here, note that $q_{1}>q_{2}\Leftrightarrow x_{1}>x_{2}$. By the direct calculation, we prove $H(\bm{X};\bm{\delta})\propto q_{1}-q_{2}$. Thus, Cor. 2 shows that after a sufficiently long time passes, $\bm{y}=\bm{y}^{*}$ holds or $H(\bm{X};\bm{\delta})$ diverges, in other words, $H(\bm{X};\bm{\delta})>0\Leftrightarrow q_{1}>q_{2}\Leftrightarrow x_{1}>x_{2}$ holds. In the former case, Y obviously reaches the Nash equilibrium strategy, so we now assume the latter case ($x_{1}>x_{2}$ holds). We can also prove that $x_{1}>x_{2}$ is equivalent to the positive definiteness of $\bm{X}^{\rm T}\bm{U}$. Because $\bm{X}^{\rm T}\bm{U}$ continues to hold after a sufficiently long time passes, $\bm{y}=\bm{y}^{*}$ holds after a further long time passes, meaning that Y’s strategy converges to its equilibrium. ∎
5 Experimental supports for theory
This section experimentally shows the global convergence to the Nash equilibrium. First, we visualize the mechanism of global convergence in matching pennies. Following Thm. 2, $H(\bm{X};\bm{\delta})$ increases with time, and $\bm{X}^{\rm T}\bm{U}$ becomes positive definite. After that, following Thm. 1, the distance from the Nash equilibrium $D(\bm{X},\bm{y})$ decreases with time. Next, we also demonstrate global convergence for several examples of more advanced games than matching pennies. We have observed the global convergence in all zero-sum games as far as we experimented, even though we could not prove it in general zero-sum games. All the simulations in this section are performed following the Runge-Kutta fourth-order method of Eqs. (5) and (6) with a sufficiently small step size of $5\times 10^{-2}$.
Figure 3: A. Trajectories of $q_{1}$ and $q_{2}$. The rainbow contour plot indicates the value of $q_{1}-q_{2}$. All the trajectories monotonically increase $q_{1}-q_{2}$ with time and converge in the area of $q_{1}>q_{2}$ in their final states. B. Trajectories of the learning dynamics. The black broken line corresponds to the region of Nash equilibria, $\bm{x}^{{\rm st}}=\bm{y}=(1/2,1/2)$. Each colored line shows a trajectory of the learning dynamics. First, the circle markers show the initial states. Following the blue lines, the trajectories diverge from the Nash equilibria ($D(\bm{X},\bm{y})$ increases with time). However, the trajectories stop to diverge and switch to converge to the Nash equilibria ($D(\bm{X},\bm{y})$ decreases), following the red lines. The star markers are the final states and correspond to one of the Nash equilibria.
5.1 Visualization of global convergence mechanism
Let us see that learning dynamics in the matching pennies converge to the equilibrium following the mechanism we provided. To analyze the matching pennies in which both the players have two actions, we use the notation of $(x_{1|j},x_{2|j})=:(x_{j},1-x_{j})$, $q_{j}:=\log x_{j}-\log(1-x_{j})$, and $(y_{1},y_{2})=:(y,1-y)$.
Fig. 3A shows the dynamics of $q_{1}$ and $q_{2}$. Here, the colors indicate the contour plot for $q_{1}-q_{2}$, showing that $H(\bm{X};\bm{\delta})\propto q_{1}-q_{2}$ monotonically increases with time and thus Thm. 2 holds. Furthermore, we also see that X’s strategy reaches the region of $q_{1}>q_{2}\Leftrightarrow x_{1}>x_{2}$ after a sufficiently long time passes. In the region, $\bm{X}^{\rm T}\bm{U}$ is positive definite, and thus Thm. 1 is applicable after sufficiently long time passes.
Next, Fig. 3B plots the global behavior of the learning dynamics, which is described by the three parameters of $x_{1}$, $x_{2}$, and $y$. The gray line shows the region that corresponds to the Nash equilibrium, i.e., $\bm{x}^{{\rm st}}(\bm{X},\bm{y})=\bm{y}=(1/2,1/2)$. Furthermore, the colored lines show example trajectories of the learning dynamics. The blue part of the line shows that $D(\bm{X},\bm{y})$ increases at the beginning of the learning dynamics. This part is in the region of $x_{1}<x_{2}$, following Thm. 1. After that, the red part shows that $D(\bm{X},\bm{y})$ decreases, and the learning dynamics converge to the equilibrium. This part is in the region of $x_{1}>x_{2}$, following Thm. 1.
5.2 Global convergence in more advanced zero-sum games
The global convergence is observed in more advanced zero-sum games than the matching pennies. Fig. 4 shows three representative examples. In all the examples, we basically assume that there is no dominant (pure-)strategy, where X’s $i=\sigma(j)$-th action is advantageous for Y’s $j$-th action, and vise versa, for the permutation $\sigma$ (i.e., $\sigma(1)=2$, $\sigma(2)=3$, $\sigma(3)=4$, and $\sigma(4)=1$). In other words, the payoff matrices satisfy $u_{i\sigma(i)}=2$ and $u_{\sigma(j)j}=-2$. Three panels take different values in the other elements of their payoff matrix.
Figure 4: Global convergence in three advanced games. In all the games, we consider the rock-paper-scissors-like game extended to the four-action case, where the second, third, fourth, and first actions win the other’s first, second, third, and fourth actions, respectively. The winner receives the payoff of $2$ (the orange blocks in the matrices for the winning of X), while the loser sends the payoff of $2$ (the blue blocks). We now introduce three variants for the other blocks in the payoff matrix. A. We set each of the other blocks by random numbers in $[-1,1]$ (the gray blocks). Then, Y’s strategy converges to the unique Nash equilibrium (the red star marker) independent of its initial state (the blue circle markers). B. We set each of the other blocks by 00, where the payoff matrix degenerates. Y’s strategy converges to one of the Nash equilibria (the line consisting of the red star markers) depending on its initial state. C. Only the block for the interaction between an action is set to $-1$, and the others are 00. If so, X’s strategy converges to the unique Nash equilibrium (the orange star markers) independent of its initial state (the green circle markers). Instead, Y’s strategies do not converge.
First, Fig. 4A shows the case where the other elements of the payoff matrix are random numbers following the uniform distribution of $[-1,1]$. In this case, the payoff matrix does not degenerate. In other words, X has no two different strategies that generate equal payoffs for all Y’s strategies. In mathematics, there exists no $\bm{a}=(a_{i})_{1\leq i\leq m_{\sf X}}\in\mathbb{R}^{m_{\sf X}}$ such that $\Sigma_{i}a_{i}\bm{u}_{i}=\bm{0}$. We observe that the single Nash equilibrium exists, and Y’s strategy always converges to its equilibrium independent of X’s and Y’s initial strategies.
Second, Fig. 4B shows the case where all the other elements take 00. In this case, the payoff matrix degenerates; Indeed, $\bm{u}_{i}+\bm{u}_{i^{\prime}}=\bm{0}$ holds for $(i,i^{\prime})=(1,3)$, $(2,4)$. Continuous Nash equilibria are seen there, $\bm{x}^{*}=r_{\sf X}(0,1/2,0,1/2)+(1-r_{\sf X})(1/2,0,1/2,0)$ and $\bm{y}^{*}=r_{\sf Y}(0,1/2,0,1/2)+(1-r_{\sf Y})(1/2,0,1/2,0)$ for all $0\leq r_{\sf X}$ and $0\leq r_{\sf Y}\leq 1$. We observe that Y’s strategy converges to one of its equilibria depending on X’s and Y’s initial strategies.
Third, Fig. 4C shows the case where the other elements take 00 in principle except for $u_{11}=-1$. In this case, the payoff matrix degenerates because $\bm{u}_{i}+\bm{u}_{i^{\prime}}=\bm{0}$ holds for $(i,i^{\prime})=(2,4)$. Furthermore, the only Nash equilibrium exists on the boundary of strategy spaces, $\bm{x}^{*}=(0,1/2,0,1/2)$, and $\bm{y}^{*}=(1/2,0,1/2,0)$. As far as we experiment, when the payoff matrix degenerates and only the boundary Nash equilibrium exists, Y’s strategy does not converge to its equilibrium exceptionally. Nevertheless, we observe that X’s strategy converges to its equilibrium instead of Y’s.
6 Conclusion
This study considered the simplest situation of memory asymmetry between two players; only player X memorizes the other’s previous action, while player Y cannot. We formulated their learning dynamics based on the replicator dynamics. Although the existence of memory complicates the dynamics, we captured the global behavior of the learning dynamics by introducing two new quantities. One is the conditional-sum divergence, which is an extension of the previous divergence to the case of games of a reactive strategy. The time change of this extended divergence is characterized by the definiteness of $\bm{X}^{\rm T}\bm{U}$. The other is a family of Lyapunov functions, characterized by $\bm{U}\log\bm{X}^{\rm T}$. This is proved to monotonically increase with time. As a valid application of these two quantities, we proved the global convergence in the learning dynamics in the simplest game equipped with an interior Nash equilibrium. We further experimentally observed the global convergence in more advanced zero-sum games. It is still a conjecture whether the learning dynamics in games of a reactive strategy converge to the Nash equilibrium even in general zero-sum payoff matrices. This study provides novel and valid indicators to analyze dynamics in learning in games.
Appendix
Appendix A Proofs
A.1 Proof of Thm. 1
Proof.
The dynamics of $D(\bm{X},\bm{y})$ are calculated as
$\dot{D}(\bm{X},\bm{y})=$ $-\Sigma_{i}\Sigma_{j}x_{i}^{*}\frac{\dot{x}_{i|j}}{x_{i|j}}-\Sigma_{j}y_{j}^{*}\frac{\dot{y}_{j}}{y_{j}}$
$=$ $-\Sigma_{i}\Sigma_{j}x_{i}^{*}\left(\frac{\partial u^{{\rm st}}}{\partial x_{i|j}}-\Sigma_{i}x_{i|j}\frac{\partial u^{{\rm st}}}{\partial x_{i|j}}\right)-\Sigma_{j}y_{j}^{*}\left(\frac{\partial u^{{\rm st}}}{\partial y_{j}}-\Sigma_{j}y_{j}\frac{\partial u^{{\rm st}}}{\partial y_{j}}\right)$
$=$ $-\Sigma_{i}\Sigma_{j}x_{i}^{*}y_{j}\left(\Sigma_{j^{\prime}}u_{ij^{\prime}}y_{j^{\prime}}-\Sigma_{i}x_{i|j}\Sigma_{j^{\prime}}u_{ij^{\prime}}y_{j^{\prime}}\right)+\Sigma_{j}y_{j}^{*}\left(\Sigma_{i}u_{ij}x_{i}^{{\rm st}}-\Sigma_{j}y_{j}\Sigma_{i}u_{ij}x_{i}^{{\rm st}}\right)$
$+\Sigma_{j}y_{j}^{*}\left(\Sigma_{i}x_{i|j}\Sigma_{j^{\prime}}u_{ij^{\prime}}y_{j^{\prime}}-\Sigma_{j}y_{j}\Sigma_{i}x_{i|j}\Sigma_{j^{\prime}}u_{ij^{\prime}}y_{j^{\prime}}\right)$
$=$ $-\underbrace{\Sigma_{i}\Sigma_{j}\Sigma_{j^{\prime}}x_{i}^{*}y_{j}u_{ij^{\prime}}y_{j^{\prime}}}_{=:{\rm(A)}=u^{*}}+\underbrace{\Sigma_{i}\Sigma_{j}\Sigma_{j^{\prime}}x_{i|j}y_{j}u_{ij^{\prime}}y_{j^{\prime}}}_{=:{\rm(B)}=u^{{\rm st}}}+\underbrace{\Sigma_{i}\Sigma_{j}y_{j}^{*}u_{ij}x_{i}^{{\rm st}}}_{=:{\rm(C)}=u^{*}}-\underbrace{\Sigma_{i}\Sigma_{j}y_{j}u_{ij}x_{i}^{{\rm st}}}_{=:{\rm(C)}=u^{{\rm st}}}$
$+\underbrace{\Sigma_{i}\Sigma_{j}\Sigma_{j^{\prime}}y_{j}^{*}x_{i|j}u_{ij^{\prime}}y_{j^{\prime}}-\Sigma_{i}\Sigma_{j}\Sigma_{j^{\prime}}y_{j}x_{i|j}u_{ij^{\prime}}y_{j^{\prime}}}_{=:{\rm(E)}=-{\rm d}\bm{y}^{T}\bm{X}^{\rm T}\bm{U}{\rm d}\bm{y}}$
$=$ $-u^{*}+u^{{\rm st}}+u^{*}-u^{{\rm st}}-{\rm d}\bm{y}^{T}\bm{X}^{\rm T}\bm{U}{\rm d}\bm{y}$
$=$ $-{\rm d}\bm{y}^{T}\bm{X}^{\rm T}\bm{U}{\rm d}\bm{y}$
((=( = $:D^{\dagger}(\bm{X};{\rm d}\bm{y})).$ (17)
Here, we calculated (A)-(D) as
${\rm(A)}$ $=\Sigma_{i}\Sigma_{j}\Sigma_{j^{\prime}}x_{i}^{*}y_{j}u_{ij^{\prime}}y_{j^{\prime}}$
$=\Sigma_{j}\Sigma_{j^{\prime}}y_{j}y_{j^{\prime}}(\Sigma_{i}x_{i}^{*}u_{ij^{\prime}})$
$=\Sigma_{j}\Sigma_{j^{\prime}}y_{j}y_{j^{\prime}}u^{*}$
$=u^{*},$ (18)
${\rm(B)}$ $=\Sigma_{i}\Sigma_{j}\Sigma_{j^{\prime}}x_{i|j}y_{j}u_{ij^{\prime}}y_{j^{\prime}}$
$=\Sigma_{i}\Sigma_{j^{\prime}}(\Sigma_{j}x_{i|j}y_{j})u_{ij^{\prime}}y_{j^{\prime}}$
$=\Sigma_{i}\Sigma_{j^{\prime}}x_{i}^{{\rm st}}u_{ij^{\prime}}y_{j^{\prime}}$
$=u^{{\rm st}},$ (19)
${\rm(C)}$ $=\Sigma_{i}\Sigma_{j}y_{j}^{*}u_{ij}x_{i}^{{\rm st}}$
$=\Sigma_{i}(\Sigma_{j}y_{j}^{*}u_{ij})x_{i}^{{\rm st}}$
$=u^{*}\Sigma_{i}x_{i}^{{\rm st}}$
$=u^{*},$ (20)
${\rm(D)}$ $=\Sigma_{i}\Sigma_{j}y_{j}u_{ij}x_{i}^{{\rm st}}$
$=u^{{\rm st}},$ (21)
${\rm(E)}$ $=\Sigma_{i}\Sigma_{j}\Sigma_{j^{\prime}}y_{j}^{*}x_{i|j}u_{ij^{\prime}}y_{j^{\prime}}-\Sigma_{i}\Sigma_{j}\Sigma_{j^{\prime}}y_{j}x_{i|j}u_{ij^{\prime}}y_{j^{\prime}}$
$=-\Sigma_{i}\Sigma_{j}\Sigma_{j^{\prime}}\underbrace{(y_{j}-y_{j}^{*})}_{=:{\rm d}y_{j}}x_{i|j}u_{ij^{\prime}}\underbrace{y_{j^{\prime}}}_{={\rm d}y_{j^{\prime}}+y_{j^{\prime}}^{*}}$
$=-\Sigma_{i}\Sigma_{j}\Sigma_{j^{\prime}}{\rm d}y_{j}x_{i|j}u_{ij^{\prime}}{\rm d}y_{j^{\prime}}-\Sigma_{i}\Sigma_{j}\Sigma_{j^{\prime}}{\rm d}y_{j}x_{i|j}u_{ij^{\prime}}y_{j^{\prime}}^{*}$
$=-\Sigma_{i}\Sigma_{j}\Sigma_{j^{\prime}}{\rm d}y_{j}x_{i|j}u_{ij^{\prime}}{\rm d}y_{j^{\prime}}-u^{*}\Sigma_{i}\Sigma_{j}y_{j}x_{i|j}$
$=-\Sigma_{i}\Sigma_{j}\Sigma_{j^{\prime}}{\rm d}y_{j}x_{i|j}u_{ij^{\prime}}{\rm d}y_{j^{\prime}}$
$=-{\rm d}\bm{y}^{\rm T}\bm{X}^{\rm T}\bm{U}{\rm d}\bm{y}.$ (22)
Here, note that ${\rm d}\bm{y}=\bm{y}-\bm{y}^{*}$ is zero-sum vector. Thus, if $\bm{X}^{\rm T}\bm{U}$ is positive definite for zero-sum vectors, $D^{\dagger}(\bm{X};{\rm d}\bm{y})\leq 0$ holds independent of ${\rm d}\bm{y}$, meaning that $D(\bm{X},\bm{y})$ monotonically decreases with time. ∎
A.2 Proof of Thm. 2
Proof.
By Eqs. (5) and (8), we obtain
$\dot{H}(\bm{X},\bm{y})=$ $\Sigma_{i}\Sigma_{i^{\prime}}\Sigma_{j}\delta_{i}u_{ij}\frac{\dot{x}_{i^{\prime}|j}}{x_{i^{\prime}|j}}\delta_{i^{\prime}}$
$=$ $\Sigma_{i}\Sigma_{i^{\prime}}\Sigma_{j}\delta_{i}u_{ij}\left(\frac{\partial u^{{\rm st}}}{\partial x_{i|j}}-\Sigma_{i}x_{i|j}\frac{\partial u^{{\rm st}}}{\partial x_{i|j}}\right)\delta_{i^{\prime}}$
$=$ $\Sigma_{i}\Sigma_{i^{\prime}}\Sigma_{j}\delta_{i}u_{ij}y_{j}\delta_{i^{\prime}}(\Sigma_{j^{\prime}}u_{i^{\prime}j^{\prime}}y_{j^{\prime}}-\Sigma_{i^{\prime\prime}}x_{i^{\prime\prime}|j}\Sigma_{j^{\prime}}u_{i^{\prime\prime}j^{\prime}}y_{j^{\prime}})$
$=$ $\Sigma_{i}\Sigma_{i^{\prime}}\Sigma_{j}\Sigma_{j^{\prime}}\delta_{i}u_{ij}y_{j}\delta_{i^{\prime}}u_{i^{\prime}j^{\prime}}y_{j^{\prime}}-\Sigma_{i}\Sigma_{i^{\prime\prime}}\Sigma_{j}\Sigma_{j^{\prime}}\underbrace{(\Sigma_{i^{\prime}}\delta_{i^{\prime}})}_{=0}\delta_{i}u_{ij}y_{j}x_{i^{\prime\prime}|j}u_{i^{\prime\prime}j^{\prime}}y_{j^{\prime}}$
$=$ $\Sigma_{i}\Sigma_{i^{\prime}}\Sigma_{j}\Sigma_{j^{\prime}}\delta_{i}u_{ij}y_{j}\delta_{i^{\prime}}u_{i^{\prime}j^{\prime}}y_{j^{\prime}}$
$=$ $(\Sigma_{i}\Sigma_{j}\delta_{i}u_{ij}y_{j})^{2}$
$=$ $(\Sigma_{i}\Sigma_{j}\delta_{i}u_{ij}{\rm d}y_{j}+\underbrace{\Sigma_{i}\Sigma_{j}\delta_{i}u_{ij}y_{j}^{*}}_{=u^{*}\Sigma_{i}\delta_{i}=0})^{2}$
$=$ $(\bm{\delta}^{\rm T}\bm{U}{\rm d}\bm{y})^{2}$
((=( = $:H^{\dagger}(\bm{y};\bm{\delta})).$ (23)
This equation shows that $H^{\dagger}(\bm{y};\bm{\delta})\geq 0$ always holds. Furthermore, $H^{\dagger}(\bm{y};\bm{\delta})=0$ for all $\bm{\delta}$ is satisfied if and only if ${\rm d}\bm{y}=\bm{0}\Leftrightarrow\bm{y}=\bm{y}^{*}$. ∎
A.3 Proof of Thm. 3
Proof.
Now, by using the special properties of two-action games, let us simplify the notation of $\bm{X}:=\{x_{i|j}\}_{i,j}$, $\bm{y}:=\{y_{j}\}_{j}$, and $\bm{\delta}:=\{\delta_{i}\}_{i}$. From the definition, $\Sigma_{i}x_{i|j}=1\Leftrightarrow x_{1|j}+x_{2|j}=1$ for all $j\in\{1,2\}$ holds, and thus we simply denote it by $x_{1|j}=:x_{j}$ and $x_{2|j}=1-x_{j}$. In a similar manner, because $y_{1}+y_{2}=1$ holds, we simply denote it by $y_{1}:=y$ and $y_{2}=1-y$. Furthermore, since $\Sigma_{i}\delta_{i}=0$ holds, we simply denote it by $\delta_{1}=:\delta$ and $\delta_{2}=-\delta$. In two-action games, we can write
$\bm{\delta}^{\rm T}$ $=\begin{pmatrix}\delta_{1}&\delta_{2}\end{pmatrix}=\begin{pmatrix}\delta&-\delta\end{pmatrix},$ (24)
$\log\bm{X}$ $=\begin{pmatrix}\log x_{1|1}&\log x_{1|2}\\ \log x_{2|1}&\log x_{2|2}\\ \end{pmatrix}=\begin{pmatrix}\log x_{1}&\log x_{2}\\ \log(1-x_{1})&\log(1-x_{2})\\ \end{pmatrix}.$ (25)
Using these notations, we calculate $H(\bm{X};\bm{\delta})$ as
$H(\bm{X};\bm{\delta})$ $=\bm{\delta}^{\rm T}\bm{U}\log\bm{X}^{\rm T}\bm{\delta}$
$=\begin{pmatrix}\delta&-\delta\end{pmatrix}\begin{pmatrix}1&-1\\ -1&1\\ \end{pmatrix}\begin{pmatrix}\log x_{1}&\log(1-x_{1})\\ \log x_{2}&\log(1-x_{2})\\ \end{pmatrix}\begin{pmatrix}\delta\\ -\delta\\ \end{pmatrix}$
$=2\delta^{2}\left(\log\frac{x_{1}}{1-x_{1}}+\log\frac{x_{2}}{1-x_{2}}\right)$
$=2\delta^{2}(q_{1}-q_{2}).$ (26)
Here, in the final line, we used $q_{j}:=\log x_{j}-\log(1-x_{j})\Leftrightarrow x_{j}=\exp(q_{j})/(1+\exp(q_{j}))$. The dynamics of $H(\bm{X};\bm{\delta})$ is also calculated as
$H^{\dagger}(\bm{y};\bm{\delta})$ $=(\bm{\delta}^{\rm T}\bm{U}{\rm d}\bm{y})^{2}$
$=\left(\begin{pmatrix}\delta&-\delta\end{pmatrix}\begin{pmatrix}1&-1\\ -1&1\\ \end{pmatrix}\begin{pmatrix}y-y^{*}\\ -y+y^{*}\\ \end{pmatrix}\right)^{2}$
$=16\delta^{2}(y-y^{*})^{2}.$ (27)
This equation shows that for $\delta\neq 0$, $H^{\dagger}(\bm{y};\bm{\delta})=0$ holds if and only if $y=y^{*}$. Thus, after sufficiently long time passes, $\bm{y}=\bm{y}^{*}$ holds or $H(\bm{X};\bm{\delta})$ diverges to infinity. If so, $H(\bm{X};\bm{\delta})>0$ is satisfied and is equivalent to $q_{1}>q_{2}$, which is also equivalent to $x_{1}>x_{2}$ from the definition. If $x_{1}>x_{2}$ holds, $D^{\dagger}(\bm{X};{\rm d}\bm{y})<0$ is satisfied because we can derive
$D^{\dagger}(\bm{X};{\rm d}\bm{y})$ $=-{\rm d}\bm{y}^{\rm T}\bm{X}^{\rm T}\bm{U}{\rm d}\bm{y}$
$=-\begin{pmatrix}y-y^{*}&-y+y^{*}\end{pmatrix}\begin{pmatrix}x_{1}&1-x_{1}\\ x_{2}&1-x_{2}\\ \end{pmatrix}\begin{pmatrix}u_{11}&u_{12}\\ u_{21}&u_{22}\\ \end{pmatrix}\begin{pmatrix}y-y^{*}\\ -y+y^{*}\\ \end{pmatrix}$
$=-(y-y^{*})^{2}(x_{1}-x_{2})(u_{11}-u_{12}-u_{21}+u_{22}).$ (28)
Since $D(\bm{X},\bm{y})$ has its lower bound, $D^{\dagger}(\bm{X};{\rm d}\bm{y})=0\Leftrightarrow y=y^{*}$ holds after sufficiently long time passes. Thus, it was proved that Y’s strategy globally converges to its Nash equilibrium strategy. ∎
Appendix B Connection with ordinary definiteness
This study considered positive definiteness with some constraint, i.e., for zero-sum vectors. However, this constrained positive definiteness is related to the ordinarily positive definiteness without constraint, introduced as follows (describing “tilde” for distinction).
Definition 2 (Positive definiteness).
A square matrix $\tilde{\bm{M}}$ is positive definite when for all vectors $\tilde{\bm{\delta}}\neq\bm{0}$, $\tilde{\bm{\delta}}\cdot(\tilde{\bm{M}}\tilde{\bm{\delta}})>0$.
Now, we show how the positive definiteness for zero-sum vectors is tied to the ordinary positive definiteness. First, freely choose and fix $\hat{k}\in\{1,\cdots,m\}$. From any $m\times m$ matrix $\bm{M}=(m_{kk^{\prime}})_{k,k^{\prime}}$, we define a $(m-1)\times(m-1)$ matrix $\tilde{\bm{M}}:=(\tilde{m}_{kk^{\prime}})_{k\neq\hat{k},k^{\prime}\neq\hat{k}}$ by
$\tilde{m}_{kk^{\prime}}$ $:=m_{kk^{\prime}}-m_{k\hat{k}}-m_{\hat{k}k^{\prime}}+m_{\hat{k}\hat{k}}.$ (29)
Then, the positive definiteness of $\bm{M}$ for zero-sum vectors is equivalent to the positive definiteness of $\tilde{\bm{M}}$ in the ordinary definition as follows.
Theorem 4 (Equivalence to positive definiteness).
If and only if $m\times m$ matrix $\bm{M}$ is positive definite for zero-sum vectors, $(m-1)\times(m-1)$ matrix $\tilde{\bm{M}}$ is positive definite.
Proof.
For any $(m-1)$-dimensional vector $\tilde{\bm{\delta}}=(\delta_{k})_{k\neq\hat{k}}$, we define a $m$-dimensional zero-sum vector $\bm{\delta}$ by
$\delta_{k}:=\begin{cases}-\Sigma_{k\neq\hat{k}}\tilde{\delta}_{k}&(k=\hat{k})\\ \tilde{\delta}_{k}&(k\neq\hat{k})\\ \end{cases}.$ (30)
Conversely, it is trivially possible to define $\tilde{\bm{\delta}}$ from any $\bm{\delta}$. Thus, there is one-to-one correspondence between the $m$-dimensional zero-sum vector $\bm{\delta}$ and the $(m-1)$-dimensional (general-sum) vector $\tilde{\bm{\delta}}$. Now, we can prove $\bm{\delta}\cdot(\bm{M}\bm{\delta})=\tilde{\bm{\delta}}\cdot(\tilde{\bm{M}}\tilde{\bm{\delta}})$ as follows.
$\bm{\delta}\cdot(\bm{M}\bm{\delta})$ $=\sum_{k}\sum_{k^{\prime}}\delta_{k}m_{kk^{\prime}}\delta_{k^{\prime}}$
$=\sum_{k\neq\hat{k}}\sum_{k^{\prime}\neq\hat{k}}\delta_{k}m_{kk^{\prime}}\delta_{k^{\prime}}+\sum_{k\neq\hat{k}}\delta_{k}m_{k\hat{k}}\delta_{\hat{k}}+\sum_{k^{\prime}\neq\hat{k}}\delta_{\hat{k}}m_{\hat{k}k^{\prime}}\delta_{k^{\prime}}+\delta_{\hat{k}}m_{\hat{k}\hat{k}}\delta_{\hat{k}}$
$=\sum_{k\neq\hat{k}}\sum_{k^{\prime}\neq\hat{k}}\tilde{\delta}_{k}m_{kk^{\prime}}\tilde{\delta}_{k^{\prime}}-\sum_{k\neq\hat{k}}\sum_{k^{\prime}\neq\hat{k}}\tilde{\delta}_{k}m_{k\hat{k}}\tilde{\delta}_{k^{\prime}}-\sum_{k\neq\hat{k}}\sum_{k^{\prime}\neq\hat{k}}\tilde{\delta}_{k}m_{\hat{k}k^{\prime}}\tilde{\delta}_{k^{\prime}}+\sum_{k\neq\hat{k}}\sum_{k^{\prime}\neq\hat{k}}\tilde{\delta}_{k}m_{\hat{k}\hat{k}}\tilde{\delta}_{k^{\prime}}$
$=\sum_{k\neq\hat{k}}\sum_{k^{\prime}\neq\hat{k}}\tilde{\delta}_{k}(m_{kk^{\prime}}-m_{k\hat{k}}-m_{\hat{k}k^{\prime}}+m_{\hat{k}\hat{k}})\tilde{\delta}_{k^{\prime}}$
$=\sum_{k\neq\hat{k}}\sum_{k^{\prime}\neq\hat{k}}\tilde{\delta}_{k}\tilde{m}_{kk^{\prime}}\tilde{\delta}_{k^{\prime}}$
$=\tilde{\bm{\delta}}\cdot(\tilde{\bm{M}}\tilde{\bm{\delta}}).$ (31)
Here, in the third equal sign, we used Eq. (30). Thus, if and only if $\bm{M}$ is positive definite for zero-sum vectors (i.e., $\bm{\delta}\cdot(\bm{M}\bm{\delta})>0$), $\tilde{\bm{M}}$ is positive definite (i.e., $\tilde{\bm{\delta}}\cdot(\tilde{\bm{M}}\tilde{\bm{\delta}})>0$). ∎
Appendix C Connection to the classical divergence
This section is dedicated to understanding the connection between our conditional-sum divergence $D(\bm{X},\bm{y})$ and the classical divergence $D_{\rm c}(\bm{x},\bm{y})$. Here, recall that X’s reactive strategy consists of multiple vectors $\bm{X}=(\bm{x}_{j})_{1\leq j\leq m_{\sf Y}}$, where $\bm{x}_{j}$ is the vector of the probability distribution of its action choice when the other chose $j$-th action in the previous round. On the other hand, in the classical divergence, X’s mixed strategy $\bm{x}$ is the single vector of the probability distribution of its action choice independent of the other’s previous choice. Thus, if the reactive strategy of $\bm{X}$ satisfies $\bm{x}_{j}=\bm{x}$ for all $j$, it corresponds to the mixed strategy of $\bm{x}$. Then, $\dot{D}(\bm{X},\bm{y})=\dot{D}_{\rm c}(\bm{x},\bm{y})$ holds, as the following theorem shows.
Theorem 5 (Connection with the classical total divergence).
When $\bm{X}$ satisfies $\bm{x}_{j}=\bm{x}$ for all $j$, $\dot{D}(\bm{X},\bm{y})=\dot{D}_{\rm c}(\bm{x},\bm{y})=0$ holds.
Proof.
Let us prove that $\dot{D}=\Sigma_{j}\dot{D}_{\rm KL}(\bm{x}^{*}\|\bm{x}_{j})+\dot{D}_{\rm KL}(\bm{y}^{*}\|\bm{y})$ and $\dot{D}_{\rm c}=\dot{D}_{\rm KL}(\bm{x}^{*}\|\bm{x})+\dot{D}_{\rm KL}(\bm{y}^{*}\|\bm{y})$ are equal. First, the time evolution of Y’s divergence, i.e., $\dot{D}_{\rm KL}(\bm{y}^{*}\|\bm{y})$, is trivially equal between these two equations. Thus, we prove $\Sigma_{j}\dot{D}_{\rm KL}(\bm{x}^{*}\|\bm{x}_{j})=\dot{D}_{\rm KL}(\bm{x}^{*}\|\bm{x})$ below.
$\Sigma_{j}\dot{D}_{\rm KL}(\bm{x}^{*}\|\bm{x}_{j})$ $=-\Sigma_{j}\Sigma_{i}x_{i}^{*}\frac{\dot{x}_{i|j}}{x_{i|j}}$
$=-\Sigma_{j}\Sigma_{i}x_{i}^{*}y_{j}(\Sigma_{j^{\prime}}u_{ij^{\prime}}y_{j^{\prime}}-\Sigma_{i}\underbrace{x_{i|j}}_{=x_{i}}\Sigma_{j^{\prime}}u_{ij^{\prime}}y_{j^{\prime}})$
$=-\Sigma_{i}x_{i}^{*}(\Sigma_{j^{\prime}}u_{ij^{\prime}}y_{j^{\prime}}-\Sigma_{i}x_{i}\Sigma_{j^{\prime}}u_{ij^{\prime}}y_{j^{\prime}})$
$=\dot{D}_{\rm KL}(\bm{x}^{*}\|\bm{x}).$ (32)
In the final equal sign, we used the replicator dynamics of X’s strategy for games without memories as
$\dot{x}_{i}$ $=x_{i}\left(\frac{{\rm d}u_{{\rm st}}(\bm{x},\bm{y})}{{\rm d}x_{i}}-\Sigma_{i}x_{i}\frac{{\rm d}u_{{\rm st}}(\bm{x},\bm{y})}{{\rm d}x_{i}}\right)$
$=x_{i}(\Sigma_{j}u_{ij}y_{j}-\Sigma_{i}x_{i}\Sigma_{j}u_{ij}y_{j})$ (33)
Thus, we proved $\dot{D}=\dot{D}_{\rm c}$. ∎
In this proof, we proved $\dot{D}(\bm{X},\bm{y})=\dot{D}_{\rm c}(\bm{x},\bm{y})$ separately for each divergence of X’s strategy and Y’s strategy.
Furthermore, note that $\dot{D}(\bm{X},\bm{y})=\dot{D}_{\rm c}(\bm{x},\bm{y})$ holds only for a moment. Because the learning dynamics give $\dot{\bm{x}}_{j}\neq\dot{\bm{x}}$, $\bm{x}_{j}$ gradually differs from $\bm{x}$. In conclusion, $\dot{D}(\bm{X},\bm{y})=\dot{D}_{\rm c}(\bm{x},\bm{y})$ becomes unsatisfied immediately. (However, $\dot{D}_{\rm c}(\bm{x},\bm{y})=0$ continues to hold.)
Appendix D Computational environment
The simulations presented in this paper were conducted using the following computational environment.
  • Operating System: macOS Monterey (version 12.4)
  • Programming Language: Python 3.11.3
  • Processor: Apple M1 Pro (10 cores)
  • Memory: 32 GB

### v2

@@
 Global Behavior of Learning Dynamics in Zero-Sum Games with Memory Asymmetry
-Abstract
+Abstract.
 This study examines the global behavior of dynamics in learning in games between two players, X and Y.
 We consider the simplest situation for memory asymmetry between two players: X memorizes the other Y’s previous action and uses reactive strategies, while Y has no memory.
-Although this memory complicates the learning dynamics, we discover two novel quantities that characterize the global behavior of such complex dynamics.
+Although this memory complicates their learning dynamics, we characterize the global behavior of such complex dynamics by discovering and analyzing two novel quantities.
 One is an extended Kullback-Leibler divergence from the Nash equilibrium, a well-known conserved quantity from previous studies.
 The other is a family of Lyapunov functions of X’s reactive strategy.
-These two quantities capture the global behavior in which X’s strategy becomes more exploitative, and the exploited Y’s strategy converges to the Nash equilibrium.
-Indeed, we theoretically prove that Y’s strategy globally converges to the Nash equilibrium in the simplest game equipped with an equilibrium in the interior of strategy spaces.
-Furthermore, our experiments also suggest that this global convergence is universal for more advanced zero-sum games than the simplest game.
-This study provides a novel characterization of the global behavior of learning in games through a couple of indicators.
-1 Introduction
-Learning in games targets how multiple agents learn their optimal strategies in the repetition of games [1].
-The set of such players’ best strategies is defined as Nash equilibrium [2], where every player has no motivation to change his/her strategy.
+One of the global behaviors we capture is that if X exploits Y, then their strategies converge to the Nash equilibrium.
+Another is that if Y’s strategy is out of equilibrium, then X becomes more exploitative with time.
+Consequently, we suggest global convergence to the Nash equilibrium from both aspects of theory and experiment.
+This study provides a novel characterization of the global behavior in learning in games through a couple of indicators.
+Key words and phrases:
+Multi-Agent Learning, Zero-Sum Game, Dynamical Systems, Lyapunov Function
+1. Introduction
+Learning in games targets how multiple agents learn their optimal strategies in the repetition of games Fudenberg and Levine (1998).
+The set of such players’ best strategies is defined as Nash equilibrium Nash Jr (1950), where every player has no motivation to change his/her strategy.
 However, this equilibrium is hard to compute in general because one’s best strategy depends on the others’ strategies.
@@
 Even when players try to learn their optimal strategies there, their strategies often cycle around the Nash equilibrium and fail to converge to the equilibrium.
-In order to understand such strange behaviors, which are unique in multi-agent learning, the dynamics of how multiple agents learn their strategies, say, learning dynamics, are frequently studied [3, 4, 5, 6].
-The representative dynamics of interest are the replicator dynamics, which is based on the evolutionary dynamics in biology [7, 8, 9, 10, 11].
-These dynamics are also known as the multiplicative weight updates (MWU) in its discrete-time version [12, 13].
-Furthermore, their connection to other representative learning dynamics, such as gradient ascent [14, 15, 16, 17] and Q-learning [18, 19, 20], should be noted.
+In order to understand such strange behaviors, which are unique in multi-agent learning, the dynamics of how multiple agents learn their strategies, say, learning dynamics, are frequently studied Tuyls et al.
+(2006); Busoniu et al.
+(2008); Tuyls and Weiss (2012); Bloembergen et al.
+(2015).
+The representative dynamics of interest are the replicator dynamics, which is based on the evolutionary dynamics in biology Taylor and Jonker (1978); Friedman (1991); Börgers and Sarin (1997); Hofbauer et al.
+(1998); Nowak and Sigmund (2004).
+These dynamics are also known as the multiplicative weight updates (MWU) in its discrete-time version Arora et al.
+(2012); Bailey and Piliouras (2018).
+Furthermore, their connection to other representative learning dynamics, such as gradient ascent Singh et al.
+(2000); Bowling and Veloso (2002); Zinkevich (2003); Bowling (2004) and Q-learning Watkins and Dayan (1992); Hussain et al.
+(2023b, a), should be noted.
 Such replicator dynamics are known to be characterized by Kullbuck-Leibler (KL) divergence, which is the distance from the Nash equilibrium to the players’ present strategies.
-This KL divergence is conserved during the learning dynamics, and the distance from the Nash equilibrium is invariant.
-Follow the Regularized Leader (FTRL) is a class of learning algorithms including the replicator dynamics and also has its conserved quantity, which is the summation of divergences for all the players [21, 22].
+This KL divergence is conserved during the learning dynamics, and the distance from the Nash equilibrium is invariant Piliouras and Shamma (2014); Piliouras et al.
+(2014).
+Follow the Regularized Leader (FTRL) is a class of learning algorithms including the replicator dynamics and also has its conserved quantity, which is the summation of divergences for all the players Mertikopoulos and Sandholm (2016); Mertikopoulos et al.
+(2018).
 To summarize, such complex learning dynamics have been discussed based on their conserved quantity.
@@
 By definition, this memory allows the agent to make more complex and intelligent decisions.
-When memory is introduced into a normal-form game, the players can achieve a wider range of strategies as the Nash equilibria (known as Folk theorem [23]).
-Furthermore, memory is also introduced into learning algorithms, such as replicator dynamics [24, 25, 26, 27] and Q-learning [28, 29, 30, 31, 32].
+When memory is introduced into a normal-form game, the players can achieve a wider range of strategies as the Nash equilibria (known as Folk theorem Fudenberg and Maskin (2009)).
+Furthermore, memory is also introduced into learning algorithms, such as replicator dynamics Fujimoto and Kaneko (2019, 2021); Fujimoto et al.
+(2023, 2024) and Q-learning Masuda and Ohtsuki (2009); Barfuss (2020); Usui and Ueda (2021); Meylahn et al.
+(2022); Ueda (2023).
 Here, since this memory causes feedback from the past, the global dynamics of such learning algorithms become more complex.
-Indeed, replicator dynamics diverge from the Nash equilibrium under symmetric memory lengths between players [26], while converging under asymmetric memory lengths [27].
+Indeed, replicator dynamics diverge from the Nash equilibrium under symmetric memory lengths between players Fujimoto et al.
+(2023), while converging under asymmetric memory lengths Fujimoto et al.
+(2024).
 Here, KL divergence is no longer useful to capture the global dynamics because it increases or decreases over time.
-The analysis of the dynamics in with-memory games is limited to the local, linearized stability analysis in the vicinity of Nash equilibrium [27].
+The analysis of the dynamics in with-memory games is limited to the local, linearized stability analysis in the vicinity of Nash equilibrium Fujimoto et al.
+(2024).
 To summarize, since memory crucially complicates learning dynamics, the global behavior of the dynamics is still unexplored.
-This study provides the first theoretical analysis of the global behavior of learning in with-memory games.
-We assume games where their memory structure is simplest and asymmetric; One side adopts a reactive strategy that can memorize the other’s previous action [33, 34, 35, 24, 36, 25], while the other has no memory.
-In order to characterize the global behavior of such with-memory games, we extend KL divergence and prove that such extended divergence increases or decreases with time depending on the reactive strategy (see Fig.
-1A).
-Furthermore, we propose a family of Lyapunov functions that characterize the dynamics of reactive strategy (see Fig.
-1B).
-These Lyapunov functions show that the with-memory side monotonically learns to exploit the no-memory side.
-As an application of these functions, we prove that the convergence from arbitrary initial strategies, i.e., global convergence, occurs in matching pennies.
-We also experimentally confirm that such global convergence is observed in general zero-sum games.
-Figure 1: A.
-Illustration of the global behavior of the conditional divergence, $D(\bm{X},\bm{y})$.
+Figure 1.
+(A) Illustration of the global behavior of the conditional divergence, $D(\bm{X},\bm{y})$.
 Three trajectories (red, black, and blue) are plotted with the Nash equilibrium (the black star marker).
-The horizontal and vertical axes show X’s strategy ($x_{1}^{{\rm st}}$) and Y’s strategy ($y_{1}$) in matching pennies (formulated in Fig.
+The horizontal and vertical axes show X’s strategy ($x_{1}^{{\rm st}}$) and Y’s strategy ($y_{1}$) in the matching pennies game (formulated in Fig.
 2).
@@
 These three lines are plotted for the different initial strategies, i.e., $\bm{X}$ and $\bm{y}$.
-B.
-Illustration of the global behavior of the family of Lyapunov functions, $H(\bm{X};\bm{\delta})$.
+(B) Illustration of the global behavior of the family of Lyapunov functions, $H(\bm{X};\bm{\delta})$.
 The colored line shows a trajectory (from purple to red) of Lyapunov functions $H_{1}$, $H_{2}$, and $H_{3}$, each of which is $H(\bm{X};\bm{\delta})$ for some specific $\bm{\delta}$.
@@
 All of $H_{1}$, $H_{2}$, and $H_{3}$ monotonically increase with time.
-2 Preliminary
-2.1 Settings
+This study provides the first theoretical analysis of the global behavior of learning in with-memory games.
+We assume games where their memory structure is simplest and asymmetric; One side adopts a reactive strategy that can memorize the other’s previous action Nowak (1990); Ohtsuki (2004); Baek et al.
+(2016); Fujimoto and Kaneko (2019); Schmid et al.
+(2022); Fujimoto and Kaneko (2021), while the other has no memory.
+In order to characterize the global behavior of such with-memory games, we extend KL divergence and prove that such extended divergence increases or decreases with time depending on whether the reactive strategy is exploitative or not (see Fig.
+1A).
+We further propose a family of Lyapunov functions that characterize the dynamics of the reactive strategy (see Fig.
+1B).
+These Lyapunov functions show that the with-memory side monotonically learns to exploit the no-memory side.
+As an application of these functions, we suggest the convergence from arbitrary initial strategies to the equilibrium, i.e., global convergence.
+We prove global convergence in the matching pennies game.
+We also experimentally confirm that such global convergence is observed in other games equipped with various types of equilibrium.
+2. Preliminary
+2.1. Settings
 First, we formulate two-player normal-form games.
@@
 Although our formulation of learning algorithms can be used for general games, this study focuses on zero-sum games.
-Figure 2: Illustration of games between reactive and zero-memory strategies.
+Figure 2.
+Illustration of games between reactive and zero-memory strategies.
 The area surrounded by the magenta dotted line shows the normal-form game.
@@
 Furthermore, in zero-sum games, Y receives $-u_{ij}$.
-Especially in the matching pennies, their actions of $1$ ($2$) correspond to the choice of “head” (“tail”) of a coin.
+Especially in the matching pennies game, their actions of $1$ ($2$) correspond to the choice of “head” (“tail”) of a coin.
 When their choices match $i=j$, X wins, i.e., $u_{11}=u_{22}=1$ (the orange blocks).
@@
 Thus, X uses a reactive strategy and can choose its action with the conditional probability of $x_{1|j}$ and $x_{2|j}$ for Y’s previous action.
-We assume that X can use reactive strategies, i.e., can change its action choice depending on the other’s previous action.
+We assume that X use reactive strategies, i.e., can change its action choice depending on the other’s previous action.
 This reactive strategy is denoted as $\bm{X}:=(x_{i|j})_{1\leq i\leq m_{\sf X},1\leq j\leq m_{\sf Y}}\in\prod_{1\leq j\leq m_{\sf Y}}\Delta^{m_{\sf X}-1}$, a matrix composed of $m_{\sf Y}$ vectors each of which are an element of a $m_{\sf X}-1$-dimensional simplex.
@@
 Thus, $\Sigma_{j}y_{j}=1$ should be satisfied.
-2.2 Stationary state and expected payoff
+2.2. Stationary State and Expected Payoff
 We now discuss the stationary state and expected payoff of repeated games.
@@
 Last, X’s expected payoff is given by $u^{{\rm st}}(\bm{x}^{{\rm st}},\bm{y}):=\Sigma_{i}\Sigma_{j}u_{ij}p_{ij}^{{\rm st}}=\Sigma_{i}\Sigma_{j}u_{ij}x_{i}^{{\rm st}}y_{j}$.
-2.3 Nash equilibrium
+2.3. Nash Equilibrium
 We now define the Nash equilibrium in the normal-form game.
@@
 From the definition, $u^{{\rm st}}(\bm{x},\bm{y})$ is the linear function for $\bm{x}$ and $\bm{y}$, and the Nash equilibrium condition is characterized by the gradient of such expected payoffs as
-$\begin{cases}\partial u^{{\rm st}}/\partial x_{i}=\Sigma_{j}u_{ij}y_{j}^{*}=C&(x_{i}^{*}>0)\\ \partial u^{{\rm st}}/\partial x_{i}=\Sigma_{j}u_{ij}y_{j}^{*}<C&(x_{i}^{*}=0)\\ \end{cases},$ (2)
-$\begin{cases}\partial u^{{\rm st}}/\partial y_{j}=\Sigma_{j}u_{ij}x_{i}^{*}=C&(y_{i}^{*}>0)\\ \partial u^{{\rm st}}/\partial y_{i}=\Sigma_{j}u_{ij}x_{i}^{*}>C&(y_{i}^{*}=0)\\ \end{cases}.$ (3)
+$\begin{cases}\partial u^{{\rm st}}/\partial x_{i}=\Sigma_{j}u_{ij}y_{j}^{*}=C&(x_{i}^{*}>0)\\ \partial u^{{\rm st}}/\partial x_{i}=\Sigma_{j}u_{ij}y_{j}^{*}\leq C&(x_{i}^{*}=0)\\ \end{cases},$ (2)
+$\begin{cases}\partial u^{{\rm st}}/\partial y_{j}=\Sigma_{i}u_{ij}x_{i}^{*}=C&(y_{j}^{*}>0)\\ \partial u^{{\rm st}}/\partial y_{j}=\Sigma_{i}u_{ij}x_{i}^{*}\geq C&(y_{j}^{*}=0)\\ \end{cases}.$ (3)
 From these conditions, for all $i$ and $j$ such that $x_{i}^{*}>0$ and $y_{j}^{*}>0$ hold, respectively, we obtain
@@
 This is the special property in zero-sum games.
-2.4 Learning algorithm: replicator dynamics
+2.4. Learning Algorithm: Replicator Dynamics
 Let us define the replicator dynamics as a representative learning algorithm. X’s and Y’s replicator dynamics are formulated as
-$\dot{x}_{i|j}$ $=x_{i|j}\left(\frac{{\rm d}u^{{\rm st}}}{{\rm d}x_{i|j}}-\Sigma_{i}x_{i|j}\frac{{\rm d}u^{{\rm st}}}{{\rm d}x_{i|j}}\right),$ (5)
+$\dot{x}_{i|j}$ $=+x_{i|j}\left(\frac{{\rm d}u^{{\rm st}}}{{\rm d}x_{i|j}}-\Sigma_{i}x_{i|j}\frac{{\rm d}u^{{\rm st}}}{{\rm d}x_{i|j}}\right),$ (5)
 $\dot{y}_{j}$ $=-y_{j}\left(\frac{{\rm d}u^{{\rm st}}}{{\rm d}y_{j}}-\Sigma_{j}y_{j}\frac{{\rm d}u^{{\rm st}}}{{\rm d}y_{j}}\right).$ (6)
-Here, following the theorems in [26], X’s and Y’s replicator dynamics include the gradient for the expected payoff $u^{{\rm st}}$.
+Here, following the theorems in Fujimoto et al.
+(2023), X’s and Y’s replicator dynamics include the gradient for the expected payoff $u^{{\rm st}}$.
 Thus, the update of X’s strategy increases its payoff $u^{{\rm st}}$, while that of Y’s strategy decreases the other’s payoff $u^{{\rm st}}$.
-3 Novel quantities characterizing learning dynamics
+We discuss learning based on the replicator dynamics throughout this study, but we can extend all the following results to another typical learning algorithm, the gradient descent-ascent (see Appendix D for detailed discussion).
+3. Theory on Learning Dynamics
 This section analyzes the dynamics of Eqs.
@@
 (5) and (6): An extended KL divergence and a family of Lyapunov functions.
-3.1 Polynomial expressions of learning dynamics
+3.1. Polynomial Expressions of Learning
 First, the gradient terms in Eqs. (5) and (6) are computed as
@@
 Notably, however, these equations are polynomial expressions with $\bm{X}$ and $\bm{y}$.
-Such polynomial expressions cannot be seen in the games of other memory lengths [26, 27] but are special in the games between reactive and no memory strategies.
-3.2 Positive definiteness for zero-sum vectors
+Such polynomial expressions cannot be seen in the games of other memory lengths Fujimoto et al.
+(2023, 2024) but are special in the games between reactive and no memory strategies.
+3.2. Positive Definiteness for Zero-Sum Vectors
 Next, let us introduce a definiteness of matrices.
@@
 The positive definiteness for zero-sum vectors connects with an ordinary positive definiteness by a simple transformation of a matrix (see Appendix B for details).
-3.3 Extended Kullback-Leibler divergence
+3.3. Extended Kullback-Leibler Divergence
 The first quantity is an extended version of divergence.
@@
 Theorem 2 (Monotonic decrease of $D$ for positive definite $\bm{X}^{\rm T}\bm{U}$).
-If $\bm{X}^{\rm T}\bm{U}$ is positive definite for zero-sum vector, $D^{\dagger}(\bm{X};{\rm d}\bm{y}):=\dot{D}(\bm{X},\bm{y})<0$ for all ${\rm d}\bm{y}\neq\bm{0}$.
+If $\bm{X}^{\rm T}\bm{U}$ is positive definite for zero-sum vector, $D^{\dagger}(\bm{X};{\rm d}\bm{y}):=\dot{D}(\bm{X},\bm{y})<0$ for all ${\rm d}\bm{y}:=\bm{y}-\bm{y}^{*}\neq\bm{0}$.
 Proof Sketch.
@@
 $\dot{D}(\bm{X},\bm{y})=-{\rm d}\bm{y}^{\rm T}\bm{X}^{\rm T}\bm{U}{\rm d}\bm{y}\ (=:D^{\dagger}(\bm{X};{\rm d}\bm{y})),$ (14)
-where we defined ${\rm d}\bm{y}:=\bm{y}-\bm{y}^{*}$, which means the difference from Y’s equilibrium strategy and is a zero-sum vector.
+which means the difference from Y’s equilibrium strategy and is a zero-sum vector.
 Thus, when $\bm{X}^{\rm T}\bm{U}$ is positive definite for zero-sum vectors, ${\rm d}\bm{y}^{\rm T}\bm{X}^{\rm T}\bm{U}{\rm d}\bm{y}$ is always positive, leading to $D^{\dagger}(\bm{X};{\rm d}\bm{y})<0$ for all ${\rm d}\bm{y}\neq\bm{0}$.
 ∎
-If $\bm{X}^{\rm T}\bm{U}$ are always positive definite for zero-sum vectors, Thm. 1 gives the global behavior of learning dynamics as below.
-Corollary 1 (Convergence of Y’s strategy).
-If $\bm{X}^{\rm T}\bm{U}$ continues to be positive definite for zero-sum vectors, $\bm{y}=\bm{y}^{*}$ holds after sufficiently long time passes.
-Proof.
-This corollary is straightforwardly proved by Thm.
-1.
-If $\bm{X}^{\rm T}\bm{U}$ is positive definite and if $\bm{y}\neq\bm{y}^{*}$, $D^{\dagger}(\bm{X};{\rm d}\bm{y})<0$ holds.
-Here, $D(\bm{X},\bm{y})$ has its lower bound ($\min_{\bm{X},\bm{y}}D(\bm{X},\bm{y})=0$).
-Thus, $D^{\dagger}(\bm{X};{\rm d}\bm{y})=0\Leftrightarrow\bm{y}=\bm{y}^{*}$ should be satisfied eventually.
-∎
-3.4 Family of Lyapunov functions
+3.4. Family of Lyapunov Functions
 Furthermore, we introduce a Lyapunov function, which characterizes the learning dynamics of X’s reactive strategy.
@@
 Theorem 3 (Monotonic increase of $H$).
-For all $\bm{\delta}$ such that $\Sigma_{i}\delta_{i}=0$, $H(\bm{X};\bm{\delta})$ monotonically increases with time.
+For all $\bm{\delta}$ such that $\Sigma_{i}\delta_{i}=0$, $H^{\dagger}(\bm{y};\bm{\delta}):=\dot{H}(\bm{X};\bm{\delta})\geq 0$. The equality holds if and only if $d\bm{y}(=\bm{y}-\bm{y}^{*})=\bm{0}$.
 Proof Sketch. By using Eq. (4), we calculate
-$\dot{H}(\bm{X};\bm{\delta})=(\bm{\delta}^{\rm T}\bm{U}{\rm d}\bm{y})^{2}\ (=:H^{\dagger}(\bm{y};\bm{\delta})).$ (16)
-This means that $H^{\dagger}(\bm{y};\bm{\delta})\geq 0$ for all $\bm{\delta}$. ∎
+$\dot{H}(\bm{X};\bm{\delta})=|\bm{\delta}^{\rm T}\bm{U}{\rm d}\bm{y}|^{2}\ (=:H^{\dagger}(\bm{y};\bm{\delta})).$ (16)
+This means that $H^{\dagger}(\bm{y};\bm{\delta})\geq 0$ for all $\bm{\delta}$. If we substitute $\bm{\delta}=\bm{x}-\bm{x}^{*}$ for some $\bm{x}\in\Delta^{m_{\sf X}-1}$, we obtain
+$H^{\dagger}(\bm{y};\bm{x}-\bm{x}^{*})=|u^{{\rm st}}(\bm{x},\bm{y})-u^{*}|^{2}.$ (17)
+For any $\bm{y}\neq\bm{y}^{*}\Leftrightarrow{\rm d}\bm{y}\neq\bm{0}$, there is $\bm{x}$ such that $u^{{\rm st}}(\bm{x},\bm{y})-u^{*}\neq 0$.
+Thus, ${\rm d}\bm{y}=\bm{0}$ is equivalent to $H^{\dagger}(\bm{y};\bm{\delta})=0$ for all $\bm{\delta}$.
+∎
 Let us interpret the function of $H(\bm{X};\bm{\delta})$.
-For simplicity, we consider a simple class of zero-sum vectors, $\bm{\delta}=\bm{e}_{i}-\bm{e}_{i^{\prime}}$ for some $i$ and $i^{\prime}\neq i$, where $\bm{e}_{i}$ denotes the unit vector for $i$-th element.
-Then, we calculate the function as $H(\bm{X};\bm{e}_{i}-\bm{e}_{i^{\prime}})=\bm{u}_{i}\cdot\log\bm{x}_{i}-\bm{u}_{i}\cdot\log\bm{x}_{i^{\prime}}-\bm{u}_{i^{\prime}}\cdot\log\bm{x}_{i}+\bm{u}_{i^{\prime}}\cdot\log\bm{x}_{i^{\prime}}$.
-The monotonic increase of $H(\bm{X};\bm{e}_{i}-\bm{e}_{i^{\prime}})$ shows that $\log\bm{x}_{i}$ becomes correlated to $\bm{u}_{i}$, and $\log\bm{x}_{i^{\prime}}$ to $\bm{u}_{i^{\prime}}$.
-Here, $\bm{x}_{i}=(x_{i|j})_{j}$ represents the vector of the probabilities to choose $i$-th action in the condition that the other chose $j$-th action in the previous round.
-$\bm{u}_{i}=(u_{ij})_{j}$ also represents the vector of the payoffs for action $i$.
-Thus, the correlation between $\log\bm{x}_{i}$ and $\bm{u}_{i}$ means that the larger its reward $u_{ij}$ is, the more frequently X returns its $i$-th action against the other’s $j$-th action, in other words, X becomes more exploitative.
-This theorem characterizes the state of learning dynamics after a sufficiently long time passes, as follows.
-Corollary 2 (Convergence of Y’s strategy or divergence of $H$).
-After sufficiently long time passes, $\bm{y}=\bm{y}^{*}$ holds or $H(\bm{X};\bm{\delta})$ diverges to infinity for all $\bm{\delta}\neq\bm{0}$ such that $\Sigma_{i}\delta_{i}=0$.
-Proof.
-Since $H(\bm{X};\bm{\delta})\geq 0$ always holds for all $\bm{\delta}$, either $H(\bm{X};\bm{\delta})=0$ or $H(\bm{X};\bm{\delta})>0$ holds in the final state.
-The former is equivalent to $\bm{y}=\bm{y}^{*}$, while the latter is equivalent to $H(\bm{X};\bm{\delta})\to\infty$.
-∎
-4 Global behavior of learning dynamics
-Now, Cors.
-1 and 2 are utilized to capture the global behavior of the learning dynamics.
-One of the global behaviors seen in the zero-sum games is the global convergence to the Nash equilibrium independent of X’s and Y’s initial strategies.
-First, Cor.
-1 guarantees that after a sufficiently long time passes, either Y’s strategy converges to the equilibrium or $H(\bm{X};\bm{\delta})$ diverges.
-Since the former trivially leads to such global convergence, we now consider the latter condition, i.e., $H(\bm{X};\bm{\delta})=\bm{\delta}^{\rm T}\bm{U}\log\bm{X}^{\rm T}\bm{\delta}\to\infty$.
-Here, $H(\bm{X};\bm{\delta})=\bm{\delta}^{\rm T}\bm{U}\log\bm{X}^{\rm T}\bm{\delta}\to\infty$ naively connects to $-D^{\dagger}(\bm{X},{\rm d}\bm{y})={\rm d}\bm{y}^{\rm T}\bm{X}^{\rm T}\bm{U}{\rm d}\bm{y}>0$ (explained in the next paragraph), meaning that $\bm{X}^{\rm T}\bm{U}$ is positive definite for zero-sum vectors.
-If so, Cor.
-2 becomes applicable, and either Y’s strategy converges to the Nash equilibrium.
-We explain the connection between $H(\bm{X};\bm{\delta})=\bm{\delta}^{\rm T}\bm{U}\log\bm{X}^{\rm T}$ and $-D^{\dagger}(\bm{X},{\rm d}\bm{y})={\rm d}\bm{y}^{\rm T}\bm{X}^{\rm T}\bm{U}{\rm d}\bm{y}>0$.
-Here, $H(\bm{X};\bm{\delta})$ and $-D^{\dagger}(\bm{X},{\rm d}\bm{y})$ have two commonalities.
-The first one is that $\bm{\delta}$ and ${\rm d}\bm{y}$ are both zero-sum vectors in common.
-The other one is that both $H$ and $-D^{\dagger}$ are given by the product of X’s strategy $\bm{X}$ and the payoff matrix $\bm{U}$.
-This means that the more $\bm{X}$ is correlated to $\bm{U}$, the larger both $H$ and $-D^{\dagger}$ are.
-Here, however, we also remark two differences of $H$ from $-D^{\dagger}$; $H$ considers the logarithm in X’s strategy $\bm{X}$ and the reversed order in the product of $\bm{X}$ and $\bm{U}$.
-In the following, we prove such global convergence of either player’s strategy in the matching pennies, the simplest game equipped with a full-support equilibrium.
-We also experimentally show global convergence in more complicated games than the matching pennies later.
-4.1 Example: Matching-pennies
-Let us define the matching-pennies game (see Fig.
-2 for the illustration of its payoff matrix).
-This game considers the action numbers of $m_{\sf X}=m_{\sf Y}=2$ and the payoff matrix of $\bm{U}=((u_{11},u_{12}),(u_{21},u_{22}))=((+1,-1),(-1,+1))$.
-The Nash equilibrium of this game is only $\bm{x}^{*}=\bm{y}^{*}=(1/2,1/2)$.
-This game is the simplest example of games that have a full-support equilibrium.
-In addition, it has been known that the replicator dynamics in games without memories cycle around the Nash equilibrium and cannot reach the equilibrium.
-Nevertheless, the learning dynamics in games of a reactive strategy achieve the convergence to the Nash equilibrium for Y’s strategy.
-The following theorem holds (see Appendix A.3 for its full proof).
-Theorem 3 (Global convergence in matching-pennies).
-In the matching-pennies ($\bm{U}=((+1,-1),(-1,+1))$), Y’s strategy converges to the Nash equilibrium ($\bm{y}=\bm{y}^{*}$), independent of both the players’ initial strategies.
-Proof Sketch.
-First, we define a special notation in two-action games; $(x_{1|j},x_{2|j})=:(x_{j},1-x_{j})$ for $j=1,2$.
-We also define the logit function of $x_{j}$ as $q_{j}:=\log x_{j}-\log(1-x_{j})$.
-Here, note that $q_{1}>q_{2}\Leftrightarrow x_{1}>x_{2}$.
-By the direct calculation, we prove $H(\bm{X};\bm{\delta})\propto q_{1}-q_{2}$.
-Thus, Cor.
-2 shows that after a sufficiently long time passes, $\bm{y}=\bm{y}^{*}$ holds or $H(\bm{X};\bm{\delta})$ diverges, in other words, $H(\bm{X};\bm{\delta})>0\Leftrightarrow q_{1}>q_{2}\Leftrightarrow x_{1}>x_{2}$ holds.
-In the former case, Y obviously reaches the Nash equilibrium strategy, so we now assume the latter case ($x_{1}>x_{2}$ holds).
-We can also prove that $x_{1}>x_{2}$ is equivalent to the positive definiteness of $\bm{X}^{\rm T}\bm{U}$.
-Because $\bm{X}^{\rm T}\bm{U}$ continues to hold after a sufficiently long time passes, $\bm{y}=\bm{y}^{*}$ holds after a further long time passes, meaning that Y’s strategy converges to its equilibrium.
-∎
-5 Experimental supports for theory
-This section experimentally shows the global convergence to the Nash equilibrium.
-First, we visualize the mechanism of global convergence in matching pennies.
-Following Thm.
-2, $H(\bm{X};\bm{\delta})$ increases with time, and $\bm{X}^{\rm T}\bm{U}$ becomes positive definite.
-After that, following Thm.
-1, the distance from the Nash equilibrium $D(\bm{X},\bm{y})$ decreases with time.
-Next, we also demonstrate global convergence for several examples of more advanced games than matching pennies.
-We have observed the global convergence in all zero-sum games as far as we experimented, even though we could not prove it in general zero-sum games.
-All the simulations in this section are performed following the Runge-Kutta fourth-order method of Eqs.
-(5) and (6) with a sufficiently small step size of $5\times 10^{-2}$.
-Figure 3: A.
-Trajectories of $q_{1}$ and $q_{2}$.
+First, $H(\bm{X};\bm{\delta})$ is the quadratic form of matrix $\bm{U}\log\bm{X}^{\rm T}$ for $\bm{\delta}$.
+We now focus on the meaning of $\bm{U}\log\bm{X}^{\rm T}$.
+The $i^{\prime}$, $i$ element of $\bm{U}\log\bm{X}^{\rm T}$ is given by $\bm{u}_{i^{\prime}}\cdot\log\bm{x}_{i}$, in which we denoted vector $\bm{x}_{i}:=(x_{i|j})_{j}$.
+This is the inner product of the payoff under X taking $i^{\prime}$-th action ($\bm{u}_{i^{\prime}}$) and the logarithmic strategy under X taking $i$-th action ($\log\bm{x}_{i}$).
+Thus, $\bm{U}\log\bm{X}^{\rm T}$ shows a correspondence matrix between X’s strategy and its payoff matrix.
+Since $H(\bm{X};\bm{\delta})$ is the quadratic form of the correspondence matrix, it tends to be large when the diagonal elements of the correspondence matrix are large.
+The diagonal elements, i.e., $\bm{u}_{i}\cdot\log\bm{x}_{i}$, indicate how X exploits Y for each $i$-th action.
+Thus, $H(\bm{X};\bm{\delta})$ means the degree of exploitation of Y by X.
+Figure 3.
+(A) Trajectories of $q_{1}$ and $q_{2}$.
 The rainbow contour plot indicates the value of $q_{1}-q_{2}$.
 All the trajectories monotonically increase $q_{1}-q_{2}$ with time and converge in the area of $q_{1}>q_{2}$ in their final states.
-B.
-Trajectories of the learning dynamics.
+(B) Trajectories of the learning dynamics.
 The black broken line corresponds to the region of Nash equilibria, $\bm{x}^{{\rm st}}=\bm{y}=(1/2,1/2)$.
@@
 The star markers are the final states and correspond to one of the Nash equilibria.
-5.1 Visualization of global convergence mechanism
-Let us see that learning dynamics in the matching pennies converge to the equilibrium following the mechanism we provided.
-To analyze the matching pennies in which both the players have two actions, we use the notation of $(x_{1|j},x_{2|j})=:(x_{j},1-x_{j})$, $q_{j}:=\log x_{j}-\log(1-x_{j})$, and $(y_{1},y_{2})=:(y,1-y)$.
+3.5. Global Behavior by Two Quantities
+So far, the two theorems (Thms. 2 and 3) explain how the two quantities vary with time. Let us understand the global behavior of the learning dynamics by interpreting the two quantities.
+$D$ explains increasing/decreasing of distance:
+First, recall that $D(\bm{X},\bm{y})$ means the distance from the Nash equilibrium.
+Thm.
+2 shows that the distance becomes smaller when $\bm{X}^{\rm T}\bm{U}$ is positive definite, whereas larger when $\bm{X}^{\rm T}\bm{U}$ is negative definite.
+Here, we focus on $\bm{X}^{\rm T}\bm{U}$ which determines whether the distance becomes smaller or larger.
+The $j$, $j^{\prime}$ element of $\bm{X}^{\rm T}\bm{U}$ is given by $\bm{x}_{j}\cdot\bm{u}_{j^{\prime}}$, in words, the correspondence between X’s strategy for $j$-th action ($\bm{x}_{j}$) and its payoff for $j^{\prime}$-th action ($\bm{u}_{j^{\prime}}$).
+The eigenvalues of a matrix are roughly determined by the diagonal elements $\bm{x}_{j}\cdot\bm{u}_{j}$.
+This diagonal element is larger when $x_{i|j}$ takes a larger value for larger $u_{ij}$, meaning that X exploits Y’s payoff more.
+As the simplest example, $\bm{X}^{\rm T}\bm{U}$ is positive definite when X takes $x_{i|j}=1$ for $i=\hat{i}$ such that $\hat{i}=\arg\max_{i}u_{ij}$, while $x_{i|j}=0$ for $i\neq\hat{i}$.
+To summarize, Thm.
+2 captures the global behavior where if X exploits Y, their strategies converge to the Nash equilibrium; otherwise, they diverge.
+$H$ explains the monotonic increase of exploitability:
+Next, recall that $\bm{U}\log\bm{X}^{\rm T}$ in $H(\bm{X};\bm{\delta})$ indicates a correspondence matrix between X’s strategy and its payoff matrix.
+Thus, Thm.
+3 explains that unless Y takes the equilibrium strategy, this correspondence continues to increase with time.
+We also pay attention to the change in the degree of correspondence, i.e., $H^{\dagger}(\bm{y};\bm{\delta})$.
+By substituting $\bm{\delta}=\bm{x}-\bm{x}^{*}$ for some $\bm{x}\in\Delta^{m_{\sf X}-1}$, we rewrite it as $H^{\dagger}(\bm{y};\bm{x}-\bm{x}^{*})=|u^{{\rm st}}(\bm{x},\bm{y})-u^{*}|^{2}$.
+Here, $|u^{{\rm st}}(\bm{x},\bm{y})-u^{*}|$ indicates the difference in payoff from the equilibrium, i.e., the exploitation by X to Y.
+Therefore, the correspondence between X’s strategy and its payoff matrix becomes larger according to the exploitability.
+Remark:
+We have interpreted both $\bm{X}^{\rm T}\bm{U}$ and $\bm{U}\log\bm{X}^{\rm T}$ as the degree of correspondence between X’s strategy and its payoff matrix.
+The interpretation is qualitatively true, but we remark that there are several quantitative differences between $\bm{X}^{\rm T}\bm{U}$ and $\bm{U}\log\bm{X}^{\rm T}$, such as the order of multiplication and the existence of logarithm.
+4. Application: Global Convergence
+By combining the global behaviors obtained from Thms.
+2 and  3, we expect that the global convergence to the Nash equilibrium occurs regardless of X’s and Y’s initial strategies.
+Thm.
+3 shows that as long as Y’s strategy is out of equilibrium, the correspondence between X’s strategy and its payoff matrix continues to be stronger.
+Afterward, Thm.
+2 shows that if the correspondence is sufficiently strong, Y’s strategy is induced to the Nash equilibrium.
+In the following, our theory and experiment support that such global convergence occurs.
+4.1. Example: Matching Pennies
+Let us define the matching pennies game (see Fig.
+2 for the illustration of its payoff matrix).
+This game considers the action numbers of $m_{\sf X}=m_{\sf Y}=2$ and the payoff matrix of $\bm{U}=((u_{11},u_{12}),(u_{21},u_{22}))=((+1,-1),(-1,+1))$.
+The Nash equilibrium of this game is only $\bm{x}^{*}=\bm{y}^{*}=(1/2,1/2)$.
+This game is the simplest example of a game equipped with a full-support Nash equilibrium.
+In addition, it has been known that the replicator dynamics in games without memories cycle around the Nash equilibrium and cannot reach the equilibrium.
+Nevertheless, by considering the memory asymmetry (Eqs.
+(5) and (6)), Y’s strategy succeeds in the convergence to the equilibrium, as shown in the following corollary (see Appendix A.3 for its full proof).
+For convenience, we use a special notation available in two-action games; $(x_{1|j},x_{2|j})=:(x_{j},1-x_{j})$, $q_{j}:=\log x_{j}-\log(1-x_{j})$, and $(y_{1},y_{2})=:(y,1-y)$.
+Corollary 0 (Global convergence in matching pennies).
+In the matching pennies game $\bm{U}=((+1,-1),(-1,+1))$, Y’s strategy $\bm{y}$ converges to the equilibrium $\bm{y}^{*}$, independent of both the players’ initial strategies.
+Proof Sketch.
+First, note that $q_{1}>q_{2}\Leftrightarrow x_{1}>x_{2}$.
+By the direct calculation, we prove $H(\bm{X};\bm{\delta})\propto q_{1}-q_{2}$.
+Thm.
+3 shows that as long as $\bm{y}=\bm{y}^{*}$, $H(\bm{X};\bm{\delta})$ continues to increase.
+Thus, after a sufficiently long time, $H(\bm{X};\bm{\delta})>0\Leftrightarrow q_{1}>q_{2}\Leftrightarrow x_{1}>x_{2}$ continue to hold.
+We can also prove that $x_{1}>x_{2}$ is equivalent to the positive definiteness of $\bm{X}^{\rm T}\bm{U}$.
+Thm.
+2 shows that under positive definite $\bm{X}^{\rm T}\bm{U}$, $\bm{y}$ asymptotically converges to $\bm{y}^{*}$, its equilibrium strategy.
+∎
+Our experiments visualize the mechanism of the global convergence based on Thm.
+2 and 3.
 Fig.
@@
 2 is applicable after sufficiently long time passes.
+Figure 4.
+Global convergence in the coupled matching pennies games, where the second, third, fourth, and first actions win the other’s first, second, third, and fourth actions, respectively.
+The winner receives the payoff of $2$ (the orange blocks in the matrices for the winning of X), while the loser sends the payoff of $2$ (the blue blocks).
+We now introduce three variants for the other blocks in the payoff matrix.
+(A) The case of interior equilibrium.
+We set each of the other blocks by random numbers in $[-1,1]$ (the gray blocks).
+Then, Y’s strategy converges to the unique Nash equilibrium (the red star marker) independent of its initial state (the blue circle markers).
+(B) The case of continuous equilibrium.
+We set each of the other blocks by 00, where the payoff matrix degenerates.
+Y’s strategy converges to one of the Nash equilibria (the line consisting of the red star markers) depending on its initial state.
+(C) The case of boundary equilibrium.
+Only the block for the interaction between an action is set to $-1$, and the others are 00.
+If so, X’s strategy converges to the unique Nash equilibrium (the orange star markers) independent of its initial state (the green circle markers).
+Instead, Y’s strategies do not converge.
 Next, Fig.
@@
 This part is in the region of $x_{1}<x_{2}$, following Thm.
-1.
+2.
 After that, the red part shows that $D(\bm{X},\bm{y})$ decreases, and the learning dynamics converge to the equilibrium.
 This part is in the region of $x_{1}>x_{2}$, following Thm.
-1.
-5.2 Global convergence in more advanced zero-sum games
-The global convergence is observed in more advanced zero-sum games than the matching pennies.
+2.
+4.2. Example: Coupled Matching Pennies
+We also observe the global convergence in other zero-sum games beyond the matching pennies game.
 Fig.
-4 shows three representative examples.
-In all the examples, we basically assume that there is no dominant (pure-)strategy, where X’s $i=\sigma(j)$-th action is advantageous for Y’s $j$-th action, and vise versa, for the permutation $\sigma$ (i.e., $\sigma(1)=2$, $\sigma(2)=3$, $\sigma(3)=4$, and $\sigma(4)=1$).
-In other words, the payoff matrices satisfy $u_{i\sigma(i)}=2$ and $u_{\sigma(j)j}=-2$.
-Three panels take different values in the other elements of their payoff matrix.
-Figure 4: Global convergence in three advanced games.
-In all the games, we consider the rock-paper-scissors-like game extended to the four-action case, where the second, third, fourth, and first actions win the other’s first, second, third, and fourth actions, respectively.
-The winner receives the payoff of $2$ (the orange blocks in the matrices for the winning of X), while the loser sends the payoff of $2$ (the blue blocks).
-We now introduce three variants for the other blocks in the payoff matrix.
-A.
-We set each of the other blocks by random numbers in $[-1,1]$ (the gray blocks).
-Then, Y’s strategy converges to the unique Nash equilibrium (the red star marker) independent of its initial state (the blue circle markers).
-B.
-We set each of the other blocks by 00, where the payoff matrix degenerates.
-Y’s strategy converges to one of the Nash equilibria (the line consisting of the red star markers) depending on its initial state.
-C.
-Only the block for the interaction between an action is set to $-1$, and the others are 00.
-If so, X’s strategy converges to the unique Nash equilibrium (the orange star markers) independent of its initial state (the green circle markers).
-Instead, Y’s strategies do not converge.
+4 considers three examples of “coupled” matching pennies games, where a pair of matching pennies games are coupled with some interaction.
+The game considers the action numbers of $m_{\sf X}=m_{\sf Y}=4$, and some elements of the payoff matrix is fixed as $u_{ij}=+2$ for $j=\sigma(i)$ and $u_{ij}=-2$ for $i=\sigma(j)$.
+Here, we used the permutation function $\sigma$ as $\sigma(1)=2$, $\sigma(2)=3$, $\sigma(3)=4$, and $\sigma(4)=1$.
+If we consider only X’s odd actions and Y’s even actions or its reverse, the payoff matrix corresponds to the matching pennies game.
+Interestingly, there are various types of Nash equilibrium depending on the interaction between these matching pennies games, i.e., the other elements of the payoff matrix, $u_{ij}$ for neither $j=\sigma(i)$ nor $i=\sigma(j)$.
+Indeed, Fig.
+4 shows three cases where Nash equilibrium exists (A) in the interior, (B) continuously, and (C) on the boundary.
+We remark that the global behavior obtained from Thm.
+2 and 3 are available even though the trajectories of the learning dynamics look complicated (see Fig.
+4).
+Furthermore, Panels A and B show that Y’s strategy converges to the equilibrium.
+The only exception is Panel C, but X’s strategy converges to the equilibrium instead of Y’s.
+In the following, we explain in detail this convergence for each panel.
+Interior equilibrium:
 First, Fig.
 4A shows the case where the other elements of the payoff matrix are random numbers following the uniform distribution of $[-1,1]$.
-In this case, the payoff matrix does not degenerate.
-In other words, X has no two different strategies that generate equal payoffs for all Y’s strategies.
-In mathematics, there exists no $\bm{a}=(a_{i})_{1\leq i\leq m_{\sf X}}\in\mathbb{R}^{m_{\sf X}}$ such that $\Sigma_{i}a_{i}\bm{u}_{i}=\bm{0}$.
-We observe that the single Nash equilibrium exists, and Y’s strategy always converges to its equilibrium independent of X’s and Y’s initial strategies.
+In this case, the payoff matrix is linearly independent.
+In mathematics, there exists no $\bm{a}=(a_{j})_{1\leq j\leq m_{\sf Y}}\in\mathbb{R}^{m_{\sf Y}}$ such that $\Sigma_{j}a_{j}\bm{u}_{j}=\bm{0}$ other than $\bm{a}=\bm{0}$.
+Thus, there is a single Nash equilibrium in the interior of the strategy space.
+As in the matching pennies game, we observe that Y’s strategy always converges to its equilibrium independent of X’s and Y’s initial strategies.
+Continuous equilibrium:
 Second, Fig.
 4B shows the case where all the other elements take 00.
-In this case, the payoff matrix degenerates; Indeed, $\bm{u}_{i}+\bm{u}_{i^{\prime}}=\bm{0}$ holds for $(i,i^{\prime})=(1,3)$, $(2,4)$.
-Continuous Nash equilibria are seen there, $\bm{x}^{*}=r_{\sf X}(0,1/2,0,1/2)+(1-r_{\sf X})(1/2,0,1/2,0)$ and $\bm{y}^{*}=r_{\sf Y}(0,1/2,0,1/2)+(1-r_{\sf Y})(1/2,0,1/2,0)$ for all $0\leq r_{\sf X}$ and $0\leq r_{\sf Y}\leq 1$.
-We observe that Y’s strategy converges to one of its equilibria depending on X’s and Y’s initial strategies.
+In this case, the payoff matrix is not linearly independent in two ways.
+Indeed, $\Sigma_{j}a_{j}\bm{u}_{j}=\bm{0}$ for $\bm{a}=(0,1,0,1)$ and $(1,0,1,0)$.
+Thus, Nash equilibria exist continuously as $\bm{x}^{*}=r_{\sf X}(0,1/2,0,1/2)+(1-r_{\sf X})(1/2,0,1/2,0)$ and $\bm{y}^{*}=r_{\sf Y}(0,1/2,0,1/2)+(1-r_{\sf Y})(1/2,0,1/2,0)$ for all $0\leq r_{\sf X}\leq 1$ and $0\leq r_{\sf Y}\leq 1$.
+Even in such continuous equilibria, we observe that Y’s strategy converges to one of the equilibria depending on X’s and Y’s initial strategies.
+Boundary equilibrium:
 Third, Fig.
 4C shows the case where the other elements take 00 in principle except for $u_{11}=-1$.
-In this case, the payoff matrix degenerates because $\bm{u}_{i}+\bm{u}_{i^{\prime}}=\bm{0}$ holds for $(i,i^{\prime})=(2,4)$.
-Furthermore, the only Nash equilibrium exists on the boundary of strategy spaces, $\bm{x}^{*}=(0,1/2,0,1/2)$, and $\bm{y}^{*}=(1/2,0,1/2,0)$.
-As far as we experiment, when the payoff matrix degenerates and only the boundary Nash equilibrium exists, Y’s strategy does not converge to its equilibrium exceptionally.
-Nevertheless, we observe that X’s strategy converges to its equilibrium instead of Y’s.
-6 Conclusion
+In this case, the payoff matrix is not linearly independent in one way.
+Indeed, $\Sigma_{j}a_{j}\bm{u}_{j}=\bm{0}$ for $\bm{a}=(0,1,0,1)$.
+The only Nash equilibrium exists on the boundary of strategy spaces, $\bm{x}^{*}=(0,1/2,0,1/2)$, and $\bm{y}^{*}=(1/2,0,1/2,0)$.
+As far as we observe our experiments, Y’s strategy fails to converge the equilibrium when the payoff matrix is not linearly independent and is equipped with only the boundary Nash equilibrium.
+Nevertheless, we observe that X’s strategy converges to the equilibrium instead of Y’s.
+We find no case that neither X’s nor Y’s strategy converges to the equilibrium.
+5. Conclusion
 This study considered the simplest situation of memory asymmetry between two players; only player X memorizes the other’s previous action, while player Y cannot.
@@
 Although the existence of memory complicates the dynamics, we captured the global behavior of the learning dynamics by introducing two new quantities.
-One is the conditional-sum divergence, which is an extension of the previous divergence to the case of games of a reactive strategy.
-The time change of this extended divergence is characterized by the definiteness of $\bm{X}^{\rm T}\bm{U}$.
-The other is a family of Lyapunov functions, characterized by $\bm{U}\log\bm{X}^{\rm T}$.
-This is proved to monotonically increase with time.
-As a valid application of these two quantities, we proved the global convergence in the learning dynamics in the simplest game equipped with an interior Nash equilibrium.
-We further experimentally observed the global convergence in more advanced zero-sum games.
-It is still a conjecture whether the learning dynamics in games of a reactive strategy converge to the Nash equilibrium even in general zero-sum payoff matrices.
-This study provides novel and valid indicators to analyze dynamics in learning in games.
+One is the conditional-sum divergence, which is an extension of the previous divergence to the case of reactive strategies.
+We proved that when X exploits Y, this conditional-sum divergence becomes smaller, meaning that their strategies converge to the Nash equilibrium.
+The other is a family of Lyapunov functions, meaning X’s exploitability to Y.
+We proved that these Lyapunov functions monotonically increase, meaning that X learns to exploit Y with time.
+As a valid application of the combination of these two quantities, we suggested the global convergence to the Nash equilibrium.
+Theoretically, we proved the global convergence in the matching pennies game, the simplest game equipped with an interior Nash equilibrium.
+Our experiments further support that global convergence occurs in coupled matching pennies games, which can have various types of Nash equilibrium structures, such as interior equilibrium, continuous equilibrium, and boundary equilibrium.
+It is still a conjecture whether the learning dynamics with memory asymmetry converge to the Nash equilibrium in general zero-sum games.
+This study provides novel and valid indicators to analyze dynamics in learning in games with memories.
+{acks}
+K. Ariu is supported by JSPS KAKENHI Grant No. 23K19986.
 Appendix
 Appendix A Proofs
-A.1 Proof of Thm. 1
+A.1. Proof of Thm. 2
 Proof.
@@
 $\dot{D}(\bm{X},\bm{y})=$ $-\Sigma_{i}\Sigma_{j}x_{i}^{*}\frac{\dot{x}_{i|j}}{x_{i|j}}-\Sigma_{j}y_{j}^{*}\frac{\dot{y}_{j}}{y_{j}}$
-$=$ $-\Sigma_{i}\Sigma_{j}x_{i}^{*}\left(\frac{\partial u^{{\rm st}}}{\partial x_{i|j}}-\Sigma_{i}x_{i|j}\frac{\partial u^{{\rm st}}}{\partial x_{i|j}}\right)-\Sigma_{j}y_{j}^{*}\left(\frac{\partial u^{{\rm st}}}{\partial y_{j}}-\Sigma_{j}y_{j}\frac{\partial u^{{\rm st}}}{\partial y_{j}}\right)$
+$=$ $-\Sigma_{i}\Sigma_{j}x_{i}^{*}\left(\frac{{\rm d}u^{{\rm st}}}{{\rm d}x_{i|j}}-\Sigma_{i}x_{i|j}\frac{{\rm d}u^{{\rm st}}}{{\rm d}x_{i|j}}\right)-\Sigma_{j}y_{j}^{*}\left(\frac{{\rm d}u^{{\rm st}}}{{\rm d}y_{j}}-\Sigma_{j}y_{j}\frac{{\rm d}u^{{\rm st}}}{{\rm d}y_{j}}\right)$
 $=$ $-\Sigma_{i}\Sigma_{j}x_{i}^{*}y_{j}\left(\Sigma_{j^{\prime}}u_{ij^{\prime}}y_{j^{\prime}}-\Sigma_{i}x_{i|j}\Sigma_{j^{\prime}}u_{ij^{\prime}}y_{j^{\prime}}\right)+\Sigma_{j}y_{j}^{*}\left(\Sigma_{i}u_{ij}x_{i}^{{\rm st}}-\Sigma_{j}y_{j}\Sigma_{i}u_{ij}x_{i}^{{\rm st}}\right)$
 $+\Sigma_{j}y_{j}^{*}\left(\Sigma_{i}x_{i|j}\Sigma_{j^{\prime}}u_{ij^{\prime}}y_{j^{\prime}}-\Sigma_{j}y_{j}\Sigma_{i}x_{i|j}\Sigma_{j^{\prime}}u_{ij^{\prime}}y_{j^{\prime}}\right)$
-$=$ $-\underbrace{\Sigma_{i}\Sigma_{j}\Sigma_{j^{\prime}}x_{i}^{*}y_{j}u_{ij^{\prime}}y_{j^{\prime}}}_{=:{\rm(A)}=u^{*}}+\underbrace{\Sigma_{i}\Sigma_{j}\Sigma_{j^{\prime}}x_{i|j}y_{j}u_{ij^{\prime}}y_{j^{\prime}}}_{=:{\rm(B)}=u^{{\rm st}}}+\underbrace{\Sigma_{i}\Sigma_{j}y_{j}^{*}u_{ij}x_{i}^{{\rm st}}}_{=:{\rm(C)}=u^{*}}-\underbrace{\Sigma_{i}\Sigma_{j}y_{j}u_{ij}x_{i}^{{\rm st}}}_{=:{\rm(C)}=u^{{\rm st}}}$
+$=$ $-\underbrace{\Sigma_{i}\Sigma_{j}\Sigma_{j^{\prime}}x_{i}^{*}y_{j}u_{ij^{\prime}}y_{j^{\prime}}}_{=:{\rm(A)}=u^{*}}+\underbrace{\Sigma_{i}\Sigma_{j}\Sigma_{j^{\prime}}x_{i|j}y_{j}u_{ij^{\prime}}y_{j^{\prime}}}_{=:{\rm(B)}=u^{{\rm st}}}+\underbrace{\Sigma_{i}\Sigma_{j}y_{j}^{*}u_{ij}x_{i}^{{\rm st}}}_{=:{\rm(C)}=u^{*}}-\underbrace{\Sigma_{i}\Sigma_{j}y_{j}u_{ij}x_{i}^{{\rm st}}}_{=:{\rm(D)}=u^{{\rm st}}}$
 $+\underbrace{\Sigma_{i}\Sigma_{j}\Sigma_{j^{\prime}}y_{j}^{*}x_{i|j}u_{ij^{\prime}}y_{j^{\prime}}-\Sigma_{i}\Sigma_{j}\Sigma_{j^{\prime}}y_{j}x_{i|j}u_{ij^{\prime}}y_{j^{\prime}}}_{=:{\rm(E)}=-{\rm d}\bm{y}^{T}\bm{X}^{\rm T}\bm{U}{\rm d}\bm{y}}$
@@
 ∎
-A.2 Proof of Thm. 2
+A.2. Proof of Thm. 3
 Proof.
@@
 $\dot{H}(\bm{X},\bm{y})=$ $\Sigma_{i}\Sigma_{i^{\prime}}\Sigma_{j}\delta_{i}u_{ij}\frac{\dot{x}_{i^{\prime}|j}}{x_{i^{\prime}|j}}\delta_{i^{\prime}}$
-$=$ $\Sigma_{i}\Sigma_{i^{\prime}}\Sigma_{j}\delta_{i}u_{ij}\left(\frac{\partial u^{{\rm st}}}{\partial x_{i|j}}-\Sigma_{i}x_{i|j}\frac{\partial u^{{\rm st}}}{\partial x_{i|j}}\right)\delta_{i^{\prime}}$
-$=$ $\Sigma_{i}\Sigma_{i^{\prime}}\Sigma_{j}\delta_{i}u_{ij}y_{j}\delta_{i^{\prime}}(\Sigma_{j^{\prime}}u_{i^{\prime}j^{\prime}}y_{j^{\prime}}-\Sigma_{i^{\prime\prime}}x_{i^{\prime\prime}|j}\Sigma_{j^{\prime}}u_{i^{\prime\prime}j^{\prime}}y_{j^{\prime}})$
-$=$ $\Sigma_{i}\Sigma_{i^{\prime}}\Sigma_{j}\Sigma_{j^{\prime}}\delta_{i}u_{ij}y_{j}\delta_{i^{\prime}}u_{i^{\prime}j^{\prime}}y_{j^{\prime}}-\Sigma_{i}\Sigma_{i^{\prime\prime}}\Sigma_{j}\Sigma_{j^{\prime}}\underbrace{(\Sigma_{i^{\prime}}\delta_{i^{\prime}})}_{=0}\delta_{i}u_{ij}y_{j}x_{i^{\prime\prime}|j}u_{i^{\prime\prime}j^{\prime}}y_{j^{\prime}}$
+$=$ $\Sigma_{i}\Sigma_{i^{\prime}}\Sigma_{j}\delta_{i}u_{ij}\left(\frac{{\rm d}u^{{\rm st}}}{{\rm d}x_{i^{\prime}|j}}-\Sigma_{i^{\prime\prime}}x_{i^{\prime\prime}|j}\frac{{\rm d}u^{{\rm st}}}{{\rm d}x_{i^{\prime\prime}|j}}\right)\delta_{i^{\prime}}$
+$=$ $\Sigma_{i}\Sigma_{i^{\prime}}\Sigma_{j}\delta_{i}u_{ij}\frac{{\rm d}u^{{\rm st}}}{{\rm d}x_{i^{\prime}|j}}\delta_{i^{\prime}}-\Sigma_{i}\Sigma_{j}\delta_{i}u_{ij}\Sigma_{i^{\prime\prime}}x_{i^{\prime\prime}|j}\frac{{\rm d}u^{{\rm st}}}{{\rm d}x_{i^{\prime\prime}|j}}\underbrace{\Sigma_{i^{\prime}}\delta_{i^{\prime}}}_{=0}$
+$=$ $\Sigma_{i}\Sigma_{i^{\prime}}\Sigma_{j}\delta_{i}u_{ij}\frac{{\rm d}u^{{\rm st}}}{{\rm d}x_{i^{\prime}|j}}\delta_{i^{\prime}}$
 $=$ $\Sigma_{i}\Sigma_{i^{\prime}}\Sigma_{j}\Sigma_{j^{\prime}}\delta_{i}u_{ij}y_{j}\delta_{i^{\prime}}u_{i^{\prime}j^{\prime}}y_{j^{\prime}}$
@@
 This equation shows that $H^{\dagger}(\bm{y};\bm{\delta})\geq 0$ always holds.
-Furthermore, $H^{\dagger}(\bm{y};\bm{\delta})=0$ for all $\bm{\delta}$ is satisfied if and only if ${\rm d}\bm{y}=\bm{0}\Leftrightarrow\bm{y}=\bm{y}^{*}$.
-∎
-A.3 Proof of Thm. 3
+Furthermore, let us prove that $H^{\dagger}(\bm{y};\bm{\delta})=0$ for all $\bm{\delta}$ is equivalent to ${\rm d}\bm{y}=\bm{0}\Leftrightarrow\bm{y}=\bm{y}^{*}$.
+Here, we substitute $\bm{\delta}=\bm{x}-\bm{x}^{*}$ for $\bm{x}\in\Delta^{m_{\sf X}-1}$ (but no problem occurs even if we generally consider $\bm{x}\in\mathbb{R}^{m_{\sf X}}$ such that $\Sigma_{i}x_{i}=1$), and then
+$\bm{\delta}^{\rm T}\bm{U}{\rm d}\bm{y}$ $=(\bm{x}-\bm{x}^{*})^{\rm T}\bm{U}(\bm{y}-\bm{y}^{*})$ (25)
+$=\underbrace{\bm{x}^{\rm T}\bm{U}\bm{y}}_{=u^{{\rm st}}(\bm{x},\bm{y})}-\underbrace{\bm{x}^{\rm T}\bm{U}\bm{y}^{*}}_{=u^{*}}-\underbrace{\bm{x}^{*\rm T}\bm{U}\bm{y}}_{=u^{*}}+\underbrace{\bm{x}^{*\rm T}\bm{U}\bm{y}^{*}}_{=u^{*}}$ (26)
+$=u^{{\rm st}}(\bm{x},\bm{y})-u^{*}.$ (27)
+We obtain
+$H^{\dagger}(\bm{y};\bm{x}-\bm{x}^{*})=|u^{{\rm st}}(\bm{x},\bm{y})-u^{*}|^{2}.$ (28)
+We use the proof by contradiction.
+Assume ${\rm d}\hat{\bm{y}}\neq\bm{0}\Leftrightarrow\hat{\bm{y}}\neq\bm{y}^{*}$ such that $u^{{\rm st}}(\bm{x},\hat{\bm{y}})=u^{*}$ for all $\bm{x}$, and then we obtain for $\bm{x}=\bm{x}^{*}$
+$u^{{\rm st}}(\bm{x}^{*},\hat{\bm{y}})=u^{*}=\max_{\bm{y}}u^{{\rm st}}(\bm{x}^{*},\bm{y})$ (29)
+$\Leftrightarrow\hat{\bm{y}}=\arg\max_{\bm{y}}u^{{\rm st}}(\bm{x}^{*},\bm{y})=\bm{y}^{*},$ (30)
+which contradicts $\hat{\bm{y}}\neq\bm{y}^{*}$. We proved $H^{\dagger}(\bm{y};\bm{\delta})=0$ for all $\bm{\delta}$ is equivalent to ${\rm d}\bm{y}=\bm{0}\Leftrightarrow\bm{y}=\bm{y}^{*}$. ∎
+A.3. Proof of Cor. 4
 Proof.
@@
 (However, $\dot{D}_{\rm c}(\bm{x},\bm{y})=0$ continues to hold.)
-Appendix D Computational environment
-The simulations presented in this paper were conducted using the following computational environment.
-  • Operating System: macOS Monterey (version 12.4)
-  • Programming Language: Python 3.11.3
-  • Processor: Apple M1 Pro (10 cores)
-  • Memory: 32 GB
+Appendix D Extension to other learning algorithms
+In the main text, we focus only on the replicator dynamics.
+However, our findings are also applicable to the gradient descent-ascent algorithm.
+In the interior of the strategy space, the learning dynamics of the gradient descent-ascent are described as
+$\dot{x}_{i|j}=+\left(\frac{{\rm d}u^{{\rm st}}}{{\rm d}x_{i|j}}-\Sigma_{i}\frac{1}{m_{\sf X}}\frac{{\rm d}u^{{\rm st}}}{{\rm d}x_{i|j}}\right),\quad\dot{y}_{j}=-\left(\frac{{\rm d}u^{{\rm st}}}{{\rm d}y_{j}}-\Sigma_{j}\frac{1}{m_{\sf Y}}\frac{{\rm d}u^{{\rm st}}}{{\rm d}y_{j}}\right),$ (41)
+which correspond to Eqs. (5) and (6), respectively.
+To capture the global behavior in the learning dynamics based on the gradient descent-ascent, we have to make small changes in $D(\bm{X},\bm{y})$ and $H(\bm{X};\bm{\delta})$ as
+$D(\bm{X},\bm{y})$ $:=\Sigma_{j}D_{2}(\bm{x}^{*},\bm{x}_{j})+D_{2}(\bm{y}^{*},\bm{y}),\quad D_{2}(\bm{p}^{*},\bm{p}):=\frac{1}{2}\|\bm{p}-\bm{p}^{*}\|^{2},$ (42)
+$H(\bm{X};\bm{\delta})$ $:=\bm{\delta}^{\rm T}\bm{U}\bm{X}^{\rm T}\bm{\delta},$ (43)
+which correspond to Eqs. (13) and (15), respectively.
+This $D(\bm{X},\bm{y})$ satisfies Thm. 2 as follows.
+$\dot{D}(\bm{X},\bm{y})$ $=\Sigma_{i}\Sigma_{j}\dot{x}_{i|j}(x_{i|j}-x_{i}^{*})+\Sigma_{j}\dot{y}_{j}(y_{j}-y_{j}^{*})$
+$=\Sigma_{i}\Sigma_{j}(x_{i|j}-x_{i}^{*})\left(\frac{{\rm d}u^{{\rm st}}}{{\rm d}x_{i|j}}+-\Sigma_{i}\frac{1}{m_{\sf X}}\frac{{\rm d}u^{{\rm st}}}{{\rm d}x_{i|j}}\right)-\Sigma_{j}(y_{j}-y_{j}^{*})\left(\frac{{\rm d}u^{{\rm st}}}{{\rm d}y_{j}}-\Sigma_{j}\frac{1}{m_{\sf Y}}\frac{{\rm d}u^{{\rm st}}}{{\rm d}y_{j}}\right)$
+$=\Sigma_{i}\Sigma_{j}(x_{i|j}-x_{i}^{*})\frac{{\rm d}u^{{\rm st}}}{{\rm d}x_{i|j}}-\Sigma_{j}(y_{j}-y_{j}^{*})\frac{{\rm d}u^{{\rm st}}}{{\rm d}y_{j}}$
+$=\underbrace{\Sigma_{i}\Sigma_{j}(x_{i|j}-x_{i}^{*})y_{j}\Sigma_{j^{\prime}}u_{ij^{\prime}}y_{j^{\prime}}}_{=({\rm B})-({\rm A})=u^{{\rm st}}-u^{*}}-\underbrace{\Sigma_{j}(y_{j}-y_{j}^{*})\Sigma_{i}u_{ij}x_{i}^{{\rm st}}}_{=:({\rm D})-({\rm C})=u^{{\rm st}}-u^{*}}-\underbrace{\Sigma_{j}(y_{j}-y_{j}^{*})\Sigma_{i}x_{i|j}\Sigma_{j^{\prime}}u_{ij^{\prime}}y_{j^{\prime}}}_{=({\rm E})={\rm d}\bm{y}^{T}\bm{X}^{\rm T}\bm{U}{\rm d}\bm{y}}$
+$=-{\rm d}\bm{y}^{T}\bm{X}^{\rm T}\bm{U}{\rm d}\bm{y}$
+$=D^{\dagger}(\bm{X};{\rm d}\bm{y}).$ (44)
+Here, we used the terms $({\rm A})$-$({\rm E})$ in Appendix A.1.
+Furthermore, this $H(\bm{X};\bm{\delta})$ also satisfies Thm. 3 as follows.
+$\dot{H}(\bm{X};\bm{\delta})$ $=\Sigma_{i}\Sigma_{i^{\prime}}\Sigma_{j}\delta_{i}u_{ij}\dot{x}_{i^{\prime}|j}\delta_{i^{\prime}}$
+$=\Sigma_{i}\Sigma_{i^{\prime}}\Sigma_{j}\delta_{i}u_{ij}\left(\frac{{\rm d}u^{{\rm st}}}{{\rm d}x_{i^{\prime}|j}}+-\Sigma_{i^{\prime\prime}}\frac{1}{m_{\sf X}}\frac{{\rm d}u^{{\rm st}}}{{\rm d}x_{i^{\prime\prime}|j}}\right)\delta_{i^{\prime}}$
+$=\Sigma_{i}\Sigma_{i^{\prime}}\Sigma_{j}\delta_{i}u_{ij}\frac{{\rm d}u^{{\rm st}}}{{\rm d}x_{i^{\prime}|j}}\delta_{i^{\prime}}+\Sigma_{i}\Sigma_{j}\delta_{i}u_{ij}\Sigma_{i^{\prime\prime}}\frac{1}{m_{\sf X}}\frac{{\rm d}u^{{\rm st}}}{{\rm d}x_{i^{\prime\prime}|j}}\underbrace{\Sigma_{i^{\prime}}\delta_{i^{\prime}}}_{=0}$
+$=\Sigma_{i}\Sigma_{i^{\prime}}\Sigma_{j}\delta_{i}u_{ij}\frac{{\rm d}u^{{\rm st}}}{{\rm d}x_{i^{\prime}|j}}\delta_{i^{\prime}}$
+$=H^{\dagger}(\bm{y};\bm{\delta}).$ (45)
+In the final line, we used the same calculation as that of Appendix A.2.
+Both the replicator dynamics and the gradient descent-ascent are included in the Follow the Regularized Leaders.
+A promising future study is to capture the global behavior in learning dynamics based on general Follow the Regularized Leaders algorithms with memories.